### v1

# One-loop QCD corrections to SSA in unweighted Drell-Yan processes

Abstract
We study one-loop QCD corrections to the single transverse spin asymmetry in Drell-Yan process. The invariant mass of virtual photon and angular distributions of final lepton in Collins-Soper frame are measured. Especially, the transverse momentum of virtual photon is integrated out. Collinear twist-3 factorization formalism is adopted for the asymmetry. We use Feynman gauge in this work. To eliminate dependent twist-3 distribution functions, equation of motion for quark is used. We find that the soft divergence from the hard pole contribution in real corrections cannot be cancelled by corresponding divergences from virtual corrections. After collinear subtraction, the hard coefficient still contains soft divergence. Thus we conclude that collinear twist-3 factorization does not hold for this asymmetry at one-loop level.
I Introduction
The collinear twist-3 factorization for single transverse spin asymmetry (SSA) in pion hadronic production has been proposed for many yearsEfremov and Teryaev (1982, 1985); Qiu and Sterman (1991, 1999). This factorization formalism can also be used to describe SSAs in direct photon productionQiu and Sterman (1992), semi-inclusive deeply inelastic scattering(SIDIS)Eguchi et al. (2006, 2007); Ji et al. (2006a) and Drell-Yan process(DY)Ji et al. (2006b, c) etc. However, a proof for the factorization is still missing for these processes. Different from twist-2 factorization, there are few one-loop corrections at twist-3 level in literatures for the SSA. So far, QCD corrections to several weighted SSAs in DY and SIDIS are calculated to one-loop levelVogelsang and Yuan (2009); Kang et al. (2013); Dai et al. (2015); Yoshida (2016); Chen et al. (2017, 2018). In these weighted SSAs, the transverse momentum $q_{\perp}$ for virtual photon in DY or $P_{h\perp}$ for final detected hadron in SIDIS is integrated out. It is found the twist-3 factorization really holds. These are non-trivial checks for the factorization. However, it does not imply the unweighted SSAs can also be factorized at the same order of $\alpha_{s}$. At tree level, the asymmetry for the angular distribution of final lepton in DY with $q_{\perp}$ integrated out was shown to be nonzero, please see Ma and Zhang (2015) and reference therein. In Ma and Zhang (2015), the tree level tensor structure of the hadronic tensor is made clear (see eq.(52) in the following), which has two parts: one part is proportional to the derivative of $\delta^{2}(q_{\perp})$ and the other part is proportional to $\delta^{2}(q_{\perp})$ itself. The former is called derived part and the latter is non-derivative part. Because the weight is proportional to $q_{\perp}$, the weighted SSA receives virtual correction only from derivative part. It has been illustrated in Chen et al. (2017) that this virtual correction is the same as the correction to the usual quark form factor, which is a twist-2 quantity. The direct calculation in Vogelsang and Yuan (2009) confirms this. If there is no weight, both derivative part and non-derivative part of the hadronic tensor can contribute to SSA. Thus, the examine of factorization based on weighted SSA is incomplete.
In this work, we study the one-loop correction to unweighted SSA in DY with $q_{\perp}$ integrated. Explicitly, we calculate the angular distribution of final lepton. To avoid possible ambiguity for soft-gluon-pole contribution, we take Feynman gauge in this work. The method of diagram expansionQiu and Sterman (1999); Eguchi et al. (2007) is adopted here. The troubles for the expansion are mainly two aspects: one is there are many dependent twist-3 distribution functions. Some of these functions contain the bad component of quark field. The other is the gauge invariant distribution functions contain gluon field strength tensor and gauge links, but it seems impossible to recover these two quantities completely. In order to solve these two problems, we use at most one longitudinal gluon($G^{+}$) to do collinear expansion, and then use equation of motion for quark field to eliminate the bad component. After these treatments we get five independent quark-gluon-quark or quark-quark correlation functions. Three of them can be identified to $q_{\partial}$, $T_{F}$ and $T_{\Delta}$. Remaining two correlation functions are dangerous. To preserve QCD gauge invariance, the hard coefficients before these two functions must be zero. Really, our calculation confirms this. This indicates our expansion scheme preserves QCD gauge invariance. QED gauge invariance for the hadronic tensor is also checked at one-loop level. In this work, we consider the contribution proportional to $\bar{q}\otimes T_{F}$. That is, for unpolarized hadron, only the contribution from twist-2 anti-quark distribution function $\bar{q}(x)$ is considered and for polarized hadron, only twist-3 quark-gluon-quark distribution function $T_{F}$ is considered. The hard coefficients from virtual and real corrections are calculated explicitly, however, it is found after collinear subtraction the final hard coefficient still contains a divergence. Since our expansion scheme preserve both QCD and QED gauge invariance, we think the divergence indicates that collinear twist-3 factorization does not hold for the SSA we consider here. Very recently, Rein et al. (2025a, b) give the one-loop correction to a single spin asymmetry $A_{UT}$ for $E_{h}d^{3}\sigma/d^{3}P_{h}$ in lepton-hadron scattering $l+p(s_{\perp})\rightarrow h+X$. The final lepton is undetected. Please see eq.(1) of Rein et al. (2025a) for the illustration of the asymmetry. No breaking of the factorization is found there. This quantity can be studied with our method and we will study this process in near future.
The structure of this paper is as follows: Sec.2 is the kinematics for lepton angular distributions in Drell-Yan process; Sec.3 is the definition of all involved twist-3 distributions and their relations resulting from equation of motion of quark field. Our expansion formalism is also presented in this section; Sec.4 contains tree level results; Sec.5 contains one-loop virtual corrections; Sec.6 contains real corrections; Sec.7 is for the renormalization of twist-3 distribution function and collinear subtraction. The final hard coefficient is also given, which contains a divergence mentioned above; In Sec.8, we give an analysis based on eikonal approximation for the hard pole contribution. It is indicated that the uncancelled divergence is a soft divergence; Sec.9 is our summary.
II Kinematics
The polarized Drell-Yan process is
$h_{A}(p_{a},s_{\perp})+h_{B}(p_{b})\rightarrow\gamma^{*}[\rightarrow e^{-}(l)e^{+}(\bar{l})]+X,$ (1)
where $h_{A}$ is a spin-$\frac{1}{2}$ hadron polarized transversely with $s_{\perp}^{\mu}$ the spin vector; $h_{B}$ is a unpolarized hadron; $X$ represents undetected hadrons; the final lepton pair(here we take electron and positron as an example) is assumed from the decay of a virtual photon, and their momenta $l^{\mu}$ and $\bar{l}^{\mu}$ are detected. As usual we introduce $q^{\mu}=l^{\mu}+\bar{l}^{\mu}$ for the virtual photon. $p_{a},p_{b}$ are momenta of hadrons. The total energy squared is $s=(p_{a}+p_{b})^{2}$.
The invariant mass squared of virtual photon $Q^{2}=q^{2}$ is a hard scale. Under Bjorken limit $Q^{2}\rightarrow\infty$ and $\tau=Q^{2}/s$ fixed, we can ignore all masses of hadrons and leptonsKodaira and Tanaka (1999). The angular distribution of final lepton we want to study is
$\frac{d\sigma}{dQ^{2}d\Omega}=\frac{\alpha_{em}^{2}}{4sQ^{4}}\int d^{n}q\delta(q^{2}-Q^{2})L_{\mu\nu}W^{\mu\nu}.$ (2)
$\Omega$ is the solid angle of final lepton with momentum $l^{\mu}$, defined in Collins-Soper(CS) frameCollins and Soper (1977). In this work, we take dimensional regularization to regulate ultraviolet(UV) and infrared(IR) divergences. The dimension of $q$-integration has been set to $n=4-\epsilon$. $L^{\mu\nu}$ and $W^{\mu\nu}$ are leptonic and hadronic tensors, respectively. That is,
$L^{\mu\nu}=$ $4(l^{\mu}\bar{l}^{\nu}+l^{\nu}\bar{l}^{\mu}-\frac{1}{2}Q^{2}g^{\mu\nu})=4(-2l^{\mu}l^{\nu}+l^{\mu}q^{\nu}+l^{\nu}q^{\mu}-\frac{1}{2}Q^{2}g^{\mu\nu}),$
$W^{\mu\nu}=$ $\int\frac{d^{n}x}{(2\pi)^{n}}e^{iq\cdot x}\sum_{X}\langle h_{B},h_{A}s_{\perp}|j^{\nu}(0)|X\rangle\langle X|j^{\mu}(x)|h_{A}s_{\perp},h_{B}\rangle.$ (3)
$j^{\mu}=\bar{\psi}\gamma^{\mu}\psi$ is electro-magnetic current. $\sum_{X}$ is the phase space integration for all possible final hadrons. At parton level, the hadrons are quarks and physical gluons. In this work we use light-cone coordinates. A four-vector $a^{\mu}$ is written as $a^{\mu}=(a^{+},a^{-},a_{\perp}^{\mu})$, with $a^{\pm}=\frac{1}{\sqrt{2}}(a^{0}\pm a^{3})$. Two light-like vectors $n^{\mu},\bar{n}^{\mu}$ are introduced, so that
$a^{+}=a\cdot n,\ a^{-}=a\cdot\bar{n},\ a_{\perp}\cdot n=a_{\perp}\cdot\bar{n}=0,\ n\cdot\bar{n}=1.$ (4)
Two transverse tensors are also introduced: transverse metric $g_{\perp}^{\mu\nu}$ and transverse anti-symmetric tensor $\epsilon_{\perp}^{\mu\nu}$ as follows:
$g_{\perp}^{\mu\nu}=g^{\mu\nu}-n^{\mu}\bar{n}^{\nu}-n^{\nu}\bar{n}^{\mu},\ \epsilon_{\perp}^{\mu\nu}=\epsilon^{-+\mu\nu}=\epsilon^{\rho\tau\mu\nu}\bar{n}_{\rho}n_{\tau}.$ (5)
Note that $\epsilon^{0123}=+1$ and $\epsilon_{\perp}^{12}=+1$ in this work.
In the center of mass(CM) frame of initial hadrons, under Bjorken limit, the masses of hadrons can be ignored, so, $p_{a}^{\mu},p_{b}^{\mu}$ become light-like, that is,
$p_{a}^{\mu}=p_{a}^{+}\bar{n}^{\mu},\ p_{b}^{\mu}=p_{b}^{-}n^{\mu}.$ (6)
The solid angle $\Omega$ is defined in CS frame, which is a rest frame of lepton pair and is obtained from CM frame by two boostsCollins and Soper (1977). The first boost is along $Z$ axis so that $q^{z}=0$ after boost; the second boost is along the direction $\vec{q}_{\perp}$, so that $\vec{q}_{\perp}=0$ after boost. $l^{\mu}$ in CS frame is parameterized as
$l_{cs}^{\mu}=(l^{0},l^{1},l^{2},l^{3})=\frac{Q}{2}(1,\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta).$ (7)
We also define $l_{cs}^{\pm}=\frac{Q}{2\sqrt{2}}(1\pm\cos\theta)$ and $l_{\perp,cs}^{\mu}$ for the longitudinal and transverse components of $l^{\mu}_{cs}$, respectively. With the two Lorentz boosts done explicitly, the momentum $l^{\mu}$ in CM frame can be expressed in terms of $l_{cs}^{\mu}$ as follows:
$l^{+}=$ $\frac{q^{+}}{E_{t}}\Big{[}\frac{E_{t}}{2}+l_{cs}^{z}-\frac{1}{Q}l_{\perp,cs}\cdot q_{\perp}\Big{]},$
$l^{-}=$ $\frac{q^{-}}{E_{t}}\Big{[}\frac{E_{t}}{2}-l_{cs}^{z}-\frac{1}{Q}l_{\perp,cs}\cdot q_{\perp}\Big{]},$
$l_{\perp}^{\mu}=$ $l_{\perp,cs}^{\mu}+(\frac{E_{t}}{Q}-1)\frac{l_{\perp,cs}\cdot q_{\perp}}{q_{\perp}^{2}}q_{\perp}^{\mu}+\frac{1}{2}q_{\perp}^{\mu}.$ (8)
As a convention, for quantities defined in CM frame, the subscription “cm” is suppressed. The dot product is defined for four-vector, that is, $a_{\perp}\cdot b_{\perp}=-\vec{a}_{\perp}\cdot\vec{b}_{\perp}$ and especially $a_{\perp}^{2}=-\vec{a}_{\perp}\cdot\vec{a}_{\perp}<0$. The transverse energy $E_{t}$ for virtual photon is $E_{t}=\sqrt{Q^{2}-q_{\perp}^{2}}\geq Q$. Above representation of lepton momentum in CM frame is crucial for our calculation. In the following we do all calculations in CM frame. The spin vector $s_{\perp}^{\mu}$ is perpendicular to $p_{a}$ in CM frame. If we let $s_{\perp}^{\mu}=(0^{+},0^{-},1,0)|s_{\perp}|$, i.e., $\vec{s}_{\perp}$ defines X-axis in CM frame, then, $\tilde{s}_{\perp}^{\mu}=\epsilon_{\perp}^{\mu\nu}s_{\perp\nu}=(0^{+},0^{-},0,1)|s_{\perp}|$, and then,
$l_{\perp,cs}\cdot\tilde{s}_{\perp}=-\frac{Q}{2}|s_{\perp}|\sin\theta\sin\phi.$ (9)
The SSA we are considering is proportional to this quantity.
In this work, we check the twist-3 factorization for two quantities $I\langle L\rangle$ and $I\langle P_{7}\rangle$, which are defined as
$I\langle L\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})L_{\mu\nu}W^{\mu\nu},$
$I\langle P_{7}\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})P_{7,\mu\nu}W^{\mu\nu},\ P_{7}^{\mu\nu}=\frac{1}{\tilde{s}_{\perp}^{2}}\Big{(}q_{\perp}^{\mu}\tilde{s}_{\perp}^{\nu}+q_{\perp}^{\nu}\tilde{s}_{\perp}^{\mu}\Big{)},$ (10)
with $\tilde{s}_{\perp}^{\mu}=\epsilon_{\perp}^{\mu\nu}s_{\perp,\nu}$. $I\langle L\rangle$ is proportional to the differential cross section listed above, i.e.,
$\frac{d\sigma}{dQ^{2}d\Omega}=\frac{\alpha_{em}^{2}}{4sQ^{4}}I\langle L\rangle.$ (11)
$I\langle P_{7}\rangle$ on the other hand is proportional to the weighted cross section studied in Chen et al. (2017). Since the weighted cross section has been shown to be factorized, $I\langle P_{7}\rangle$ here is used as a check of our calculation.
III Twist-3 distribution functions
In this work we study the contribution from chiral-even twist-3 distribution functions. At twist-3 level, there are only two independent chiral-even quark-gluon-quark correlation functions,
$\tilde{s}_{\perp}^{\rho}T_{F}(x_{1},x_{2})=$ $g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{4\pi}e^{-i\xi^{-}k^{+}-i\xi_{1}^{-}k_{1}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\mathcal{L}_{n}^{\dagger}(\xi^{-})\gamma^{+}G_{\perp}^{+\rho}(\xi^{-})\mathcal{L}_{n}(\xi^{-})\mathcal{L}^{\dagger}_{n}(\xi_{1}^{-})\psi(\xi_{1}^{-})|ps\rangle,$
$s_{\perp}^{\rho}T_{\Delta}(x_{1},x_{2})=$ $g_{s}\int\frac{d\xi^{-}d\xi_{1}^{-}}{4\pi}e^{-i\xi^{-}k^{+}-i\xi_{1}^{-}k_{1}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\mathcal{L}_{n}^{\dagger}(\xi^{-})(-i)\gamma^{+}\gamma_{5}G_{\perp}^{+\rho}(\xi^{-})\mathcal{L}_{n}(\xi^{-})\mathcal{L}^{\dagger}_{n}(\xi_{1}^{-})\psi(\xi_{1}^{-})|ps\rangle,$ (12)
where the gauge link $\mathcal{L}_{n}$ ensures that the two distributions are gauge invariant. The definition of gauge link is
$\mathcal{L}_{n}(\xi^{-})=Pe^{-ig_{s}\int_{-\infty}^{\xi^{-}}d\lambda^{-}G^{+}(\lambda^{-})},\ G^{+}=G^{+}_{a}T_{a},$ (13)
with $T_{a}$ the generator of fundamental representation of $SU(N_{c})$. $P$ is path ordering operator:
$PG^{+}(\lambda_{1}^{-})G^{+}(\lambda_{2}^{-})=\theta(\lambda_{1}^{-}-\lambda_{2}^{-})G^{+}(\lambda_{1}^{-})G^{+}(\lambda_{2}^{-})+\theta(\lambda_{2}^{-}-\lambda_{1}^{-})G^{+}(\lambda_{2}^{-})G^{+}(\lambda_{1}^{-}).$ (14)
In Drell-Yan process, the gauge link points to -$\infty$. For parton momenta, throughout the paper we use following notations
$k^{+}=xp_{a}^{+},\ k_{1}^{+}=x_{1}p_{a}^{+},\ k_{2}^{+}=x_{2}p_{a}^{+},\ k_{b}^{-}=x_{b}p_{b}^{-},$ (15)
and $k_{2}=k+k_{1}$, $x_{2}=x+x_{1}$. $k_{b}$ is the momentum of anti-quark from unpolarized hadron.
In addition to these three-point distributions, there are three two-point distributions as follows
$q_{T}(x)s_{\perp}^{\rho}=$ $p^{+}\int\frac{d\xi^{-}}{4\pi}e^{i\xi^{-}xp^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\gamma_{\perp}^{\rho}\gamma_{5}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})|ps\rangle,$
$-iq^{\prime}_{\partial}(x)\tilde{s}_{\perp}^{\rho}=$ $\int\frac{d\xi^{-}}{4\pi}e^{i\xi^{-}xp^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\gamma^{+}\partial_{\perp}^{\rho}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})|ps\rangle,$
$-iq_{\partial}(x)\tilde{s}_{\perp}^{\rho}=$ $\int\frac{d\xi^{-}}{4\pi}e^{i\xi^{-}xp^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)\gamma^{+}\gamma_{5}\partial_{\perp}^{\rho}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})|ps\rangle.$ (16)
However, they are not independent due to
$\frac{1}{2\pi}\int dx_{1}P\frac{1}{x_{1}-x_{2}}\Big{[}T_{F}(x_{1},x_{2})+T_{\Delta}(x_{1},x_{2})\Big{]}=-x_{2}q_{T}(x_{2})+q_{\partial}(x_{2}),\ T_{F}(x,x)=2q^{\prime}_{\partial}(x).$ (17)
These distribution functions and relations between them can be found for example in Boer et al. (2003); Zhou et al. (2010); Kanazawa et al. (2016). $q_{T}$ and $q^{\prime}_{\partial}$ can be eliminated. Another kind of twist-3 distributions involving covariant derivative can be expressed by the distribution functions introduced aboveZhou et al. (2010). So, we expect the factorized cross section can be expressed by $T_{F},T_{\Delta}$ and $q_{\partial}$.
It is difficult to recover $T_{F},T_{\Delta}$ and $q_{\partial}$ in practical calculation, because of the gluon field strength tensor and gauge links. We do not try to recover the complete gluon field strength tensor and gauge links in this calculation. Instead, we first do collinear expansion using the matrix elements containing at most one $G^{+}$, which are
$\langle ps|\bar{\psi}\tilde{\Gamma}\psi|ps\rangle,\ \langle ps|\bar{\psi}\tilde{\Gamma}G^{+}\psi|ps\rangle,\ \langle ps|\bar{\psi}\Gamma\partial_{\perp}^{\rho}\psi|ps\rangle,\ \langle ps|\bar{\psi}\Gamma(\partial_{\perp}^{\rho}G^{+})\psi|ps\rangle,\ \langle ps|\bar{\psi}\Gamma G^{+}\partial_{\perp}^{\rho}\psi|ps\rangle,$ (18)
with $\tilde{\Gamma}=\gamma_{\perp}^{\rho},\gamma_{\perp}^{\rho}\gamma_{5}$, $\Gamma=\gamma^{+},\gamma^{+}\gamma_{5}$. Then, we try to eliminate the dependent matrix elements by using equation of motion(EOM) and parity and time reversal(PT) symmetries. Such an expansion scheme can be understood because the operators in $q_{\partial}$, $T_{F}$ and $T_{\Delta}$ can always be expanded into $\bar{\psi}_{+}$, $\psi_{+}$, $G^{+}$ and $G_{\perp}$, where $\bar{\psi}_{+}$ and $\psi_{+}$ are good components of fermion field( see eq.(23)). Then, if the factorization for $q_{\partial}$, $T_{F}$ and $T_{\Delta}$ is right, preserving the expansion to a certain power of $G^{+}$ is also right, in the sense that the coefficient functions of the resulting matrix elements are finite. In our scheme, we keep the expansion to $O(G^{+})$ and $(G_{\perp})^{0}$. Details of the expansion is given below.
We introduce following three types of correlation functions. The first type does not contain $G^{+}$,
$\int\frac{d\xi^{-}}{2\pi}e^{ik^{+}\xi^{-}}\langle Ps|\bar{\psi}_{j}(0)\psi_{i}(\xi^{-})|Ps\rangle=$ $\frac{1}{4N_{c}}\Big{[}\gamma^{-}2q(x)+\gamma_{5}\gamma_{\perp}^{\rho}s_{\perp\rho}M^{(0)}_{\gamma_{\perp}\gamma_{5}}+\gamma_{\perp}^{\rho}\tilde{s}_{\perp\rho}M^{(0)}_{\gamma_{\perp}}\Big{]}$,
$\int\frac{d\xi^{-}}{2\pi}e^{ik^{+}\xi^{-}}\langle Ps|\bar{\psi}_{j}(0)\partial_{\perp}^{\rho}\psi_{i}(\xi^{-})|Ps\rangle=$ $\frac{1}{4N_{c}}\Big{[}i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}+\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(0)}_{\gamma^{+},\partial_{\perp}\psi}\Big{]}$. (19)
The superscript $(0)$ implies there is no gluon. The subscript represents the gamma matrix in the correlation function and derivatives. From PT symmetry, it can be shown
$M^{(0)}_{\gamma_{\perp}}=M^{(0)}_{\gamma^{+},\partial\psi}=0.$ (20)
So, there are only two twist-3 two-point correlation functions, which are related to $\gamma_{5}$. $q(x)$ is the usual unpolarized twist-2 parton distribution function(PDF) for quark.
The second type contains one $G^{+}$ but no $\partial_{\perp}$,
$g_{s}\int\frac{d\xi^{-}}{2\pi}\int\frac{d\xi_{1}^{-}}{2\pi}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)G_{a}^{+}(\xi^{-})\psi_{i}(\xi_{1}^{-})|ps\rangle=\frac{T_{a}}{4N_{c}C_{F}}\Big{[}i\gamma_{\perp}^{\rho}\tilde{s}_{\perp\rho}M^{(1)}_{\gamma_{\perp}}+\gamma_{5}\gamma_{\perp}^{\rho}s_{\perp\rho}M^{(1)}_{\gamma_{\perp}\gamma_{5}}\Big{]}$; (21)
The third type contains one $G^{+}$ and one $\partial_{\perp}$,
$g_{s}\int\frac{d\xi^{-}}{2\pi}\int\frac{d\xi_{1}^{-}}{2\pi}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)G_{a}^{+}(\xi^{-})\partial_{\perp}^{\rho}\psi_{i}(\xi_{1}^{-})|ps\rangle=$ $\frac{T_{a}}{4N_{c}C_{F}}\Big{[}\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}+i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}\Big{]}$;
$g_{s}\int\frac{d\xi^{-}}{2\pi}\int\frac{d\xi_{1}^{-}}{2\pi}e^{ik^{+}\xi^{-}+ik_{1}^{+}\xi_{1}^{-}}\langle ps|\bar{\psi}_{j}(0)[\partial_{\perp}^{\rho}G_{a}^{+}(\xi^{-})]\psi_{i}(\xi_{1}^{-})|ps\rangle=$ $\frac{T_{a}}{4N_{c}C_{F}}\Big{[}\gamma^{-}\tilde{s}_{\perp}^{\rho}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}+i\gamma_{5}\gamma^{-}s_{\perp}^{\rho}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}\Big{]}$. (22)
From PT symmetry of QCD, all of these $M^{(i)}$ are real. Note that for $M^{(1)}$, one $g_{s}$ is included in the matrix element.
For fermion field, the good and bad components are defined as
$\psi_{+}=\frac{\gamma^{-}\gamma^{+}}{2}\psi,\ \psi_{-}=\frac{\gamma^{+}\gamma^{-}}{2}\psi.$ (23)
[…]
$\psi_{-}(\xi^{-})=-\frac{1}{2}\mathcal{L}_{n}(\xi^{-})\int_{-\infty}^{\xi^{-}}d\lambda^{-}\mathcal{L}_{n}^{\dagger}(\lambda^{-})\gamma^{+}\gamma_{\perp}\cdot D_{\perp}(\lambda^{-})\psi_{+}(\lambda^{-}).$ (24)
The suppression is caused by covariant derivative $D_{\perp}^{\rho}=\partial_{\perp}^{\rho}+ig_{s}G_{\perp}^{\rho}$. This relation enables us to eliminate the bad component. After eliminating the bad component we get following useful relations between $M^{(i)}$,
$M^{(1)}_{\gamma_{\perp}\gamma_{5}}(k^{+},k_{1}^{+})\doteq$ $-\frac{1}{2}P\Big{(}\frac{1}{k_{1}^{+}}+\frac{1}{k_{2}^{+}}\Big{)}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})-\frac{1}{2}P\Big{(}\frac{1}{k_{1}^{+}}-\frac{1}{k_{2}^{+}}\Big{)}M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})$
$-\frac{1}{2k_{2}^{+}}\Big{(}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})-M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{)},$
$M^{(1)}_{\gamma_{\perp}}(k^{+},k_{1}^{+})\doteq$ $\frac{1}{2}P\Big{(}\frac{1}{k_{1}^{+}}+\frac{1}{k_{2}^{+}}\Big{)}M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+\frac{1}{2}P\Big{(}\frac{1}{k_{1}^{+}}-\frac{1}{k_{2}^{+}}\Big{)}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})$
$+\frac{1}{2k_{2}^{+}}\Big{(}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})-M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{)},$
$\ M^{(0)}_{\gamma_{\perp}\gamma_{5}}(k_{2}^{+})\doteq$ $-\frac{1}{k_{2}^{+}}\int dk^{+}P\frac{1}{k_{1}^{+}}\Big{(}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\partial_{\perp}\psi}(k^{+},k_{1}^{+})\Big{)}-\frac{1}{k_{2}^{+}}M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k_{2}^{+}).$ (25)
The $\doteq$ implies $G_{\perp}$ and higher order of $G^{+}$ such as $(G^{+})^{2}$ terms are ignored on right hand side. P means principal value, and $k_{2}^{+}=k^{+}+k_{1}^{+}$. To one-loop level, a consistent treatment of $\gamma_{5}$ is important. We adopt HVBM scheme’t Hooft and Veltman (1972); Breitenlohner and Maison (1977) in this work. In this scheme $\gamma_{5}$ is defined as a four dimensional quantity. In addition, spin vectors $s_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$ are also defined as four dimensional quantities. Thus following identities can be applied
[…]
which reduce the number of gamma matrices.
Now our calculation scheme is clear. First, we use all possible twist-3 matrix elements $M^{(i)}$ to do collinear expansion and get all corresponding hard coefficients. For hadronic tensor, the result is
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}dk_{2}^{+}\bar{q}(k_{b}^{-})\Big{[}H_{0}^{\mu\nu}M_{\gamma_{\perp}\gamma_{5}}^{(0)}(k_{2}^{+})+\tilde{H}_{0}^{\mu\nu}M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k_{2}^{+})\Big{]}$
$+\int dk_{b}^{-}dk^{+}dk_{1}^{+}\bar{q}(k_{b}^{-})\Big{[}H_{1}^{\mu\nu}M^{(1)}_{\gamma_{\perp}\gamma_{5}}(k^{+},k_{1}^{+})+H_{2}^{\mu\nu}M^{(1)}_{\gamma_{\perp}}(k^{+},k_{1}^{+})$
$+H_{3}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5},\partial\psi}(k^{+},k_{1}^{+})+H_{4}^{\mu\nu}M^{(1)}_{\gamma^{+},\partial\psi}(k^{+},k_{1}^{+})+H_{5}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5},\partial G^{+}}(k^{+},k_{1}^{+})+H_{6}^{\mu\nu}M^{(1)}_{\gamma^{+},\partial G^{+}}(k^{+},k_{1}^{+})\Big{]}.$ (27)
The hard coefficients are
[…]
$\tilde{H}^{\mu\nu}_{0}=$ $\text{Tr}\Big{[}C_{F}i\frac{\partial H^{\mu\nu}}{\partial k_{2\perp}^{\tau}}\otimes i\gamma_{5}\gamma^{-}s_{\perp}^{\tau}\otimes\gamma^{+}\Big{]},$
[…]
$H^{\mu\nu}_{2}=$ $\text{Tr}\Big{[}H^{\mu\nu}\otimes iT^{a}\not{\tilde{s}}_{\perp}\otimes\gamma^{+}\Big{]},$
$H^{\mu\nu}_{3}=$ $\text{Tr}\Big{[}i\frac{\partial H^{\mu\nu}}{\partial k_{1\perp\tau}}\otimes iT^{a}\gamma_{5}\gamma^{-}s_{\perp}^{\tau}\otimes\gamma^{+}\Big{]},$
$H^{\mu\nu}_{4}=$ $\text{Tr}\Big{[}i\frac{\partial H^{\mu\nu}}{\partial k_{1\perp\tau}}\otimes T^{a}\gamma^{-}\tilde{s}_{\perp\tau}\otimes\gamma^{+}\Big{]},$
$H^{\mu\nu}_{5}=$ $\text{Tr}\Big{[}i\frac{\partial H^{\mu\nu}}{\partial k_{\perp\tau}}\otimes iT^{a}\gamma^{5}\gamma^{-}s_{\perp}^{\tau}\otimes\gamma^{+}\Big{]},$
$H^{\mu\nu}_{6}=$ $\text{Tr}\Big{[}i\frac{\partial H^{\mu\nu}}{\partial k_{\perp\tau}}\otimes T^{a}\gamma^{-}\tilde{s}_{\perp}^{\tau}\otimes\gamma^{+}\Big{]}.$ (28)
$\otimes$ is the product in Dirac and color space.
Then we use EOM relations eq.(25) to eliminate dependent correlation functions to get
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}g_{0}^{\mu\nu}\times M^{(0)}_{\gamma^{+}\gamma_{5}\partial\psi}(k_{2}^{+})$
$+\int dk_{b}^{-}\bar{q}(x_{b})\int dk^{+}dk_{1}^{+}\Big{[}g_{1}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\partial\psi}+g_{2}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\gamma_{5}\partial\psi}+g_{3}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\partial_{\perp}G^{+}}+g_{4}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\gamma_{5}\partial_{\perp}G^{+}}\Big{]}$, (29)
with
$g_{0}=$ $-\frac{1}{k_{2}^{+}}H_{0}+\tilde{H}_{0},$
$k_{2}^{+}g_{1}=$ $-P\frac{1}{k_{1}^{+}}H_{0}-P\frac{k^{+}}{2k_{1}^{+}}H_{1}+P\frac{k_{1}^{+}+k_{2}^{+}}{2k_{1}^{+}}H_{2}+k_{2}^{+}H_{4},$
$k_{2}^{+}g_{2}=$ $-P\frac{1}{k_{1}^{+}}H_{0}-P\frac{k_{1}^{+}+k_{2}^{+}}{2k_{1}^{+}}H_{1}+P\frac{k^{+}}{2k_{1}^{+}}H_{2}+k_{2}^{+}H_{3},$
$k_{2}^{+}g_{3}=$ $\frac{1}{2}H_{1}+\frac{1}{2}H_{2}+k_{2}^{+}H_{6},$
$k_{2}^{+}g_{4}=$ $-\frac{1}{2}H_{1}-\frac{1}{2}H_{2}+k_{2}^{+}H_{5}.$ (30)
$\mu,\nu$ indices in $H_{i}$ and $g_{i}$ are suppressed. Since all $M^{(i)}$ are real, the coefficients $g_{i}$ are also real for symmetric $\mu,\nu$.
Besides EOM relations, there is one more important relation if $k^{+}=0$, i.e.,
$M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})\Big{|}_{k^{+}=0}=M^{(1)}_{\gamma^{+},\partial_{\perp}\bar{\psi}}(k^{+},k_{1}^{+})\Big{|}_{k^{+}=0}=-\frac{1}{2}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{|}_{k^{+}=0}.$ (31)
This is equivalent to the relation between $q^{\prime}_{\partial}$ and $T_{F}$, and can be derived from PT symmetry. It is possible that the coefficient $g_{1}$ contains a soft-gluon-pole(SGP) part, that is,
$g_{1}(k^{+},k_{1}^{+})=\tilde{g}_{1}(k^{+},k_{1}^{+})+\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$ (32)
where $\tilde{g}_{1}(k^{+},k_{1}^{+})$ is finite at $k^{+}=0$. If this is the case, then
$\int dk^{+}dk_{1}^{+}\Big{[}g_{1}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+g_{3}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{]}$
$=$ $\int dk^{+}dk_{1}^{+}\Big{[}\tilde{g}_{1}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+\Big{(}g_{3}(k^{+},k_{1}^{+})-\frac{1}{2}\delta(k^{+})g_{1}^{SGP}(k_{1}^{+})\Big{)}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{]}.$ (33)
On the other hand, since $q_{\partial}$ contains gauge link $\mathcal{L}_{n}$ while $M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}$ does not, we should extract some parts of $g_{1},g_{3}$ to produce the gauge link $\mathcal{L}_{n}$, if we want to write $M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}$ into $q_{\partial}$. This is not difficult if we notice that
$\int dk^{+}\frac{e^{ik^{+}(\xi^{-}-\xi_{1}^{-})}}{k^{+}+i\epsilon}G^{+}(\xi^{-})=-2\pi i\theta(\xi_{1}^{-}-\xi^{-})G^{+}(\xi^{-}),$ (34)
and
$\int dk^{+}P\frac{1}{k^{+}}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}$= $\frac{1}{2}\int dk^{+}(\frac{1}{k^{+}+i\epsilon}+\frac{1}{k^{+}-i\epsilon})M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}=\frac{1}{2}\int dk^{+}\frac{1}{k^{+}+i\epsilon}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}+c.c.$ (35)
Then,
$s_{\perp}^{\rho}2\int dk^{+}P\frac{1}{k^{+}}\Big{(}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{)}$
$=$ $s_{\perp}^{\rho}M^{(0)}_{\gamma^{+}\gamma_{5},\partial\psi}(k_{2}^{+})-\int\frac{d\xi^{-}}{2\pi}e^{i\xi^{-}k_{2}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)(-i\gamma^{+}\gamma_{5})\partial_{\perp}^{\rho}\Big{(}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})\Big{)}|ps\rangle+O((G^{+})^{2}).$ (36)
Or equivalently,
$M^{(0)}_{\gamma^{+}\gamma_{5},\partial\psi}(k_{2}^{+})-2\int dk^{+}P\frac{1}{k^{+}}\Big{(}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{)}=-2q_{\partial}(x_{2})+O((G^{+})^{2}).$ (37)
Then,
$\int dk_{2}^{+}g_{0}(k_{2}^{+})M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k_{2}^{+})+\int dk^{+}dk_{1}^{+}[g_{2}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}+g_{4}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$]
$=$ $-2\int dk_{2}^{+}g_{0}(k_{2}^{+})q_{\partial}(x_{2})$
$+\int dk^{+}dk_{1}^{+}\Big{[}\Big{(}g_{2}(k^{+},k_{1}^{+})+P\frac{2g_{0}(k_{2}^{+})}{k^{+}}\Big{)}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}+\Big{(}g_{4}(k^{+},k_{1}^{+})+P\frac{2g_{0}(k_{2}^{+})}{k^{+}}\Big{)}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}\Big{]}$. (38)
After gauge link is recovered, remaining $G^{+}$ can be viewed as a part of gluon field strength tensor. We replace $\partial_{\perp}^{\rho}G^{+}$ in $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$ and $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$ into $G_{\perp}^{+\rho}$ as done in Qiu and Sterman (1999) or
$M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\rightarrow-\frac{1}{\pi}T_{F}(x_{1},x_{2}),\ M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\rightarrow-\frac{1}{\pi}T_{\Delta}(x_{1},x_{2}),$ (39)
with $x_{2}=x+x_{1}$, $k^{+}=xp_{a}^{+}$, $k_{1}^{+}=x_{1}p_{a}^{+}$. The final formula is
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}\tilde{g}_{0}^{\mu\nu}q_{\partial}(x_{2})$
$+\int dk_{b}^{-}\bar{q}(x_{b})\int dk^{+}dk_{1}^{+}\Big{[}\tilde{g}_{1}^{\mu\nu}M^{(1)}_{\gamma^{+}\partial\psi}+\tilde{g}_{2}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5}\partial\psi}-\frac{1}{\pi}\tilde{g}_{3}^{\mu\nu}T_{F}(x_{1},x_{2})-\frac{1}{\pi}\tilde{g}_{4}^{\mu\nu}T_{\Delta}(x_{1},x_{2})\Big{]},$ (40)
with
$\tilde{g}_{0}=$ $-2g_{0},$
$\tilde{g}_{1}=$ $g_{1}-\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$
$\tilde{g}_{2}=$ $g_{2}+P\frac{2g_{0}(k_{2}^{+})}{k^{+}},$
$\tilde{g}_{3}=$ $g_{3}-\frac{1}{2}\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$
$\tilde{g}_{4}=$ $g_{4}+P\frac{2g_{0}(k_{2}^{+})}{k^{+}}.$ (41)
This is our main formula for calculation. Now it is clear that if $\tilde{g}_{1}$ or $\tilde{g}_{2}$ is nonzero, the collinear expansion we employed does not preserve QCD gauge invariance. This is an important check of our calculation.
IV Tree level result
According to our expansion scheme, at most one $G^{+}$ appears. There are only two diagrams at tree level, as shown in Fig.1. The contribution of conjugated diagram of Fig.1(b) is included but not shown.
Figure 1: Tree diagrams contributing to $W^{\mu\nu}$. The conjugated diagram of (b) is not shown, but included in the calculation.
Suppose $\vec{p}_{a}$ is along +Z+Z axis in CM frame. $p_{a}^{\mu}\simeq(p_{a}^{+},0,0_{\perp})$, $p_{b}^{\mu}\simeq(0,p_{b}^{-},0_{\perp})$. Under collinear limit, the partons connecting the hard part and hadrons are collinear, whose momenta are
$k_{2}^{\mu}=(k_{2}^{+},k_{2}^{-},k_{2\perp})\sim Q(1,\lambda^{2},\lambda),\ k_{b}^{\mu}=(k_{b}^{+},k_{b}^{-},k_{b\perp})\sim Q(\lambda^{2},1,\lambda),$ (42)
with $\lambda$ a small quantity. $k$ and $k_{1}$ are also collinear to $p_{a}$, like $k_{2}$. At twist-3 level, the hard part should be expanded to $O(\lambda)$. One can find the details about collinear expansion in Qiu and Sterman (1999); Eguchi et al. (2007); Chen et al. (2017). Because the hard coefficients $H_{i}$ contain the delta function for momentum conservation $\delta^{n}(k_{2}+k_{b}-q)$, we also need to do power expansion for it, that is,
$\delta^{n-2}(k_{2\perp}-q_{\perp})=\delta^{(n-2)}(q_{\perp})-\frac{\delta^{(n-2)}(q_{\perp})}{\partial q_{\perp}^{\rho}}k_{2\perp}^{\rho}+O(k_{2\perp}^{2}).$ (43)
$\partial\delta^{n-2}(q_{\perp})/\partial q_{\perp}^{\rho}$ gives the derivative part and $\delta^{(n-2)}(q_{\perp})$ gives the non-derivative part. Next we calculate these two parts separately.
Because final leptons are unpolarized, the leptonic tensor $L^{\mu\nu}$ is symmetric in $\mu,\nu$. So, we just need the symmetric part of the hadronic tensor. For non-derivative part, only $H_{5,6}^{\mu\nu}$ can have a symmetric part. We have
$H_{5}^{\mu\nu}|_{tree}=$ $H_{6}^{\mu\nu}|_{tree}=-4iN_{c}C_{F}\frac{1}{k_{b}\cdot k+i\epsilon}\frac{1}{p_{a}^{+}}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu}).$ (44)
We just need the real part. Because
$\frac{1}{k_{b}\cdot k+i\epsilon}=P\frac{1}{k_{b}\cdot k}-i\pi\delta(k_{b}\cdot k),$ (45)
only the delta function gives a real part. This is called soft-gluon-pole(SGP) contribution, since the delta function forces the gluon momentum $k^{+}$ to be zero. We have
$H_{5}^{\mu\nu}|_{tree}=$ $H_{6}^{\mu\nu}|_{tree}=-4N_{c}C_{F}\delta(k_{b}\cdot k)\frac{1}{p_{a}^{+}}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu}).$ (46)
However, from PT symmetry,
$M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{|}_{k^{+}=0}=0.$ (47)
So, $H_{5}$ can be ignored. We have
$8N_{c}^{2}C_{F}W^{\mu\nu}_{non-de}=$ $-4\pi N_{c}C_{F}\int dk_{b}^{-}\int dk_{2}^{+}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})\Big{[}\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial G^{+}}(0,k_{2}^{+})\frac{(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu})}{p_{a}\cdot q}\Big{]}.$ (48)
For the derivative part, $H_{0,1,2}^{\mu\nu}$ do not appear, because corresponding matrix elements contain a bad component of fermion field. The contributions are of twist-4 at least. So, only $\tilde{H}_{0}^{\mu\nu}$ and $H_{3-6}^{\mu\nu}$ may contribute. However, $\tilde{H}_{0}^{\mu\nu}$ and $H_{3,5}^{\mu\nu}$ contain one $\gamma_{5}$. After taking the trace, they are proportional to $\epsilon_{\perp}^{\mu\nu}$, which is anti-symmetric in $\mu,\nu$. Thus, $\tilde{H}_{0}^{\mu\nu}$ and $H_{3,5}^{\mu\nu}$ do not contribute. As a result, we just need to calculate $H_{4}^{\mu\nu}$ and $H_{6}^{\mu\nu}$, which are
$H_{4}^{\mu\nu}=i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial k_{1\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}],\ H_{6}^{\mu\nu}=i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial k_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}].$ (49)
From Fig.1 without conjugated diagrams, we get
$p_{a}^{+}H_{4}^{\mu\nu}|_{tree}=p_{a}^{+}H_{6}^{\mu\nu}|_{tree}=-i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}8N_{c}C_{F}\frac{1}{x+i\epsilon}g_{\perp}^{\mu\nu}.$ (50)
With conjugated diagrams taken into account, only real part contributes, which is proportional to $\delta(x)$. With the help of eq.(31), $\partial_{\perp}\psi$ can be converted to $\partial_{\perp}G^{+}$. Then, we have
$8N_{c}^{2}C_{F}W^{\mu\nu}_{de}=$ $\int dk_{b}^{-}\int dk^{+}dk_{1}^{+}\bar{q}(k_{b}^{-})\Big{(}-\frac{1}{2}H_{4}^{\mu\nu}+H_{6}^{\mu\nu}\Big{)}M^{(1)}_{\gamma^{+},\partial G^{+}}$
$=$ $-4\pi N_{c}C_{F}\int dk_{b}^{-}\int dk_{2}^{+}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial G^{+}}(0,k_{2}^{+})\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}g_{\perp}^{\mu\nu}.$ (51)
Thus, the total tree level result is
$W_{tree}^{\mu\nu}=$ $W^{\mu\nu}_{non-de}+W^{\mu\nu}_{de}$
$=$ $\frac{-\pi}{2N_{c}}\int dk_{b}^{-}\int dk_{2}^{+}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial G^{+}}(0,k_{2}^{+})\Big{[}\frac{\delta^{n-2}(q_{\perp})}{p_{a}\cdot q}(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu})+\tilde{s}_{\perp}^{\rho}\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}g_{\perp}^{\mu\nu}\Big{]}$
$=$ $\frac{1}{2N_{c}}\int\frac{dx_{b}}{x_{b}}\frac{dx_{2}}{x_{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta^{n-2}(q_{\perp})\frac{1}{p_{a}\cdot q}(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu})+\tilde{s}_{\perp}^{\rho}\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}g_{\perp}^{\mu\nu}\Big{]}.$ (52)
with
$\xi\equiv\frac{q^{+}}{p_{a}^{+}},\ \tau_{\xi}\equiv\frac{\tau}{\xi}=\frac{Q^{2}}{s\xi},\ \hat{x}_{2}\equiv\frac{\xi}{x_{2}},\ \hat{x}_{b}\equiv\frac{\tau_{\xi}}{x_{b}}.$ (53)
In the last equality we have used eq.(39) to change $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$ to $T_{F}$. Above $W^{\mu\nu}_{tree}$ agrees with known result in Ma and Zhang (2015). It has been pointed out in Ma and Zhang (2015), such a structure of $W^{\mu\nu}$ with a derivative in $\delta^{(n-2)}(q_{\perp})$ satisfies QED gauge invariance in the sense of distribution, i.e.,
$\int d^{n-2}q_{\perp}t(q_{\perp})q_{\mu}W_{tree}^{\mu\nu}=0,$ (54)
with $t(q_{\perp})$ a test function which is normal at $q_{\perp}=0$.
V Virtual corrections
In this section we present our results for one-loop virtual corrections. We first give the corrections to hadronic tensor $W^{\mu\nu}$. Same as tree level $W^{\mu\nu}$, the virtual correction contains nonderivative part and derivative part, which are calculated separately. Direct calculation of the one-loop integrals is complicated because a lot of tensor integrals are involved. A better method is to use FIRESmirnov (2008) to reduce these tensor integrals to standard scalar integrals. The reduced integrals are very simple: only standard two-point integrals remain. In the calculation, both UV and IR divergences are regulated by dimensional regularization, and we do not distinguish UV and IR divergences. UV divergences will be cancelled by counter term contributions discussed in Sec.7. After corrections to $W^{\mu\nu}$ are obtained, we give the result for $I\langle L\rangle$ and $I\langle P_{7}\rangle$.
The diagrams we consider in this part are shown in Fig.2. In order to get the real part of $W^{\mu\nu}$ in physical region, we have to make clear the analyticity of the amplitude about $s_{0}$ and $s_{1}$, where
$s_{0}=(k+k_{b})^{2}=2k\cdot k_{b},\ s_{1}=(k_{1}+k_{b})^{2}=2k_{1}\cdot k_{b}.$ (55)
We also define $s_{2}=s_{0}+s_{1}=2k_{2}\cdot k_{b}$, but do not use it to eliminate $s_{0}$ or $s_{1}$ in the amplitude. The elimination will break the analytic property about $s_{0}$ or $s_{1}$. Taking $s_{0},s_{1}$ as variables is crucial for the extraction of real part of $W^{\mu\nu}$.
Figure 2: Diagrams for the hard part of one-loop virtual correction to $W^{\mu\nu}$. The right part of (a-h) is not drawn, which is a tree level photon-quark vertex. The last diagram (i) contains both left part and right part. All conjugated diagrams are not shown, but included in the calculation.
By using Feynman parameters, it can be shown that for the diagrams in Fig.2, the hard part of $W^{\mu\nu}$ is analytic on the upper half planes of $s_{0}$ and $s_{1}$, respectively. For example, one of the scalar integrals appearing for Fig.2(f) is
$I=\int\frac{d^{n}k_{g}}{(2\pi)^{n}}\frac{1}{[(k_{b}+k_{g})^{2}+i\epsilon][(k_{b}+k+k_{g})^{2}+i\epsilon][(k_{1}-k_{g})^{2}+i\epsilon][k_{g}^{2}+i\epsilon]}.$ (56)
With Feynman parameters and momentum shift, $k_{g}$ can be integrated out. Then,
$I\sim\int_{0}^{1}\prod dx_{i}\delta(1-x_{1}-x_{2}-x_{3}-x_{4})\Delta^{\frac{n}{2}-4},$ (57)
with
$\Delta=(x_{1}k_{b}+x_{2}(k_{b}+k)-x_{3}k)^{2}-x_{2}(k_{b}+k)^{2}-i\epsilon=-x_{2}(1-x_{1}-x_{2})s_{0}-x_{3}(x_{1}+x_{2})s_{1}-i\epsilon.$ (58)
Since $0\leq x_{i}\leq 1$ and $1-x_{1}-x_{2}=x_{3}+x_{4}>0$, the integral is well defined if $s_{0}$ and $s_{1}$ have positive imaginary parts. It can be checked that all integrals appearing in Fig.2 have such a feature. So, we conclude that the hard part is analytic on the upper half planes of $s_{0},s_{1}$. Further, there are only three massive quantities in $W^{\mu\nu}$, i.e., $s_{0},s_{1},s_{0}+s_{1}$, any scalar integral $I$ can be written into following form
$I=\frac{1}{(s_{0}+i\epsilon)^{\alpha}(s_{1}+i\epsilon)^{\beta}(s_{0}+s_{1}+i\epsilon)^{\gamma}}f(s_{0},s_{1}),$ (59)
with $f(s_{0},s_{1})$ a polynomial of $s_{0},s_{1}$, and $\alpha,\beta,\gamma$ some constants depending on $\epsilon=4-n$. After this is clear, following calculation is straightforward. Reduced by FIRE, all $H_{i}$ can be expressed by three two-point integrals:
$B(s_{0})=$ $\mu^{\epsilon}\int\frac{d^{n}l}{(2\pi)^{n}}\frac{1}{[l^{2}+i\epsilon][(l+k_{b0})^{2}+i\epsilon]},$
$B(s_{1})=$ $\mu^{\epsilon}\int\frac{d^{n}l}{(2\pi)^{n}}\frac{1}{[l^{2}+i\epsilon][(l+k_{b1})^{2}+i\epsilon]},$
$B(s_{2})=$ $\mu^{\epsilon}\int\frac{d^{n}l}{(2\pi)^{n}}\frac{1}{[l^{2}+i\epsilon][(l+k_{b2})^{2}+i\epsilon]},$ (60)
with $k_{b0}=k_{b}+k$, $k_{b1}=k_{b}+k_{1}$, $k_{b2}=k_{b}+k_{2}$. Moreover, the complex conjugate of $B(s_{2})$ is denoted by $B_{c}(s_{2})$, i.e., $B_{c}(s_{2})=B^{*}(s_{2})$. The expression of $B(u)$ for a general $u$ is
$B(u)=$ $i\frac{(4\pi\mu^{2})^{\epsilon/2}(-u-i\epsilon)^{-\epsilon/2}}{16\pi^{2}}\Gamma(1+\frac{\epsilon}{2})\frac{2}{\epsilon}B(1-\frac{\epsilon}{2},1-\frac{\epsilon}{2}).$ (61)
If $u<0$, $B(u)$ is purely imaginary. Expansion in $\epsilon$ gives
$B(u)=$ $i\frac{1}{16\pi^{2}}\Big{(}\frac{4\pi\mu^{2}}{|u|}\Big{)}^{\epsilon/2}\Gamma(1+\frac{\epsilon}{2})\Big{[}\frac{2}{\epsilon}+2+(2-\frac{\pi^{2}}{12})\epsilon+\theta(u)\Big{(}-\frac{\pi^{2}\epsilon}{4}\Big{)}+i\pi(1+\epsilon)\theta(u)+O(\epsilon^{2})\Big{]}.$ (62)
Now, the real part of $B(u)$ is made explicit. The real part is finite.
We list nonderivative and derivative contributions separately in the following.
V.1 Corrections to nonderivative part of $W^{\mu\nu}$
Define $h_{i}^{\mu\nu}$ as
$H_{i}^{\mu\nu}=\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})h_{i}^{\mu\nu},\ i=0,1,\cdots,6.$ (63)
$\tilde{h}_{0}^{\mu\nu}$ is defined similarly. By definition, $q_{\perp}=0$ in $h_{i}$. There is only one transverse vector in $h_{i}$, that is $\tilde{s}_{\perp}$. So, the tensor structure for $h_{i}$ is,
$h_{i}^{\mu\nu}=$ $2g_{s}^{2}\text{Re}\Big{\{}A_{i}t_{a}^{\mu\nu}+B_{i}t_{b}^{\mu\nu}\Big{\}},\ i=0,\cdots,6;$
$t_{a}^{\mu\nu}=$ $p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu},\ t_{b}^{\mu\nu}=p_{b}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{b}^{\nu}\tilde{s}_{\perp}^{\mu}.$ (64)
Only real part is needed, as indicated by $\text{Re}\{\cdots\}$. The overall factor 2 is due to conjugated diagrams. The coefficients $A_{i},B_{i}$ are obtained directly from the diagrams in Fig.2. As stated before, they are analytic on the upper half-planes of $s_{0},s_{1}$.
For $h_{0}$, we have
$\frac{1}{N_{c}C_{F}}p_{b}^{-}A_{0}=$ $-\frac{x_{2}(\epsilon-2)C_{F}}{x_{b}}B\left(s_{2}\right),$
$\frac{1}{N_{c}C_{F}}p_{b}^{-}B_{0}=$ $-\frac{\left(\epsilon^{2}+8\right)C_{F}}{\epsilon}B\left(s_{2}\right).$ (65)
For $\tilde{h}_{0}$, we have
$\frac{1}{N_{c}C_{F}}\tilde{A}_{0}=$ $-\frac{2x_{2}(\epsilon-2)C_{F}}{s_{2}}B\left(s_{2}\right),$
$\frac{1}{N_{c}C_{F}}\tilde{B}_{0}=$ $-\frac{2x_{b}\left(\epsilon^{2}+8\right)C_{F}}{s_{2}\epsilon}B\left(s_{2}\right).$ (66)
For $h_{1}$, we have
$\frac{1}{N_{c}C_{F}}A_{1}=$ $\frac{2x_{2}}{s_{2}(x+i\epsilon)}\frac{-2+\epsilon}{\epsilon}\Big{[}xC_{A}(B(s_{0})-B(s_{2}))-\epsilon x_{2}C_{F}(B(s_{2})-B_{c}(s_{2}))\Big{]},$
$\frac{1}{N_{c}C_{F}}B_{1}=$ $\frac{2x_{b}}{s_{2}(x+i\epsilon)}\frac{1}{\epsilon}\Big{[}C_{A}(-2+\epsilon)\Big{(}x_{2}B(s_{0})-xB(s_{2})\Big{)}+C_{F}(8+\epsilon^{2})x_{2}\Big{(}B(s_{2})+B_{c}(s_{2})\Big{)}\Big{]}.$ (67)
For $h_{2}$, we have
$\frac{1}{N_{c}C_{F}}A_{2}=$ $\frac{2x_{2}}{s_{2}(x+i\epsilon)}\frac{-2+\epsilon}{\epsilon}\Big{[}C_{A}x\Big{(}B(s_{0})-B(s_{2})\Big{)}-C_{F}x_{2}\epsilon\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}\Big{]},$
$\frac{1}{N_{c}C_{F}}B_{2}=$ $\frac{2x_{b}}{s_{2}(x+i\epsilon)}\frac{1}{\epsilon}\Big{[}C_{A}(-2+\epsilon)\Big{(}xB(s_{2})-x_{2}B(s_{0})\Big{)}+C_{F}(8-4\epsilon+\epsilon^{2})x_{2}\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}\Big{]}.$ (68)
For $h_{3}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}A_{3}=$ $-\frac{2x_{2}}{s_{2}(x+i\epsilon)(x_{1}+i\epsilon)}\frac{-2+\epsilon}{\epsilon}\Big{[}C_{A}x\Big{(}B(s_{0})-B(s_{2})\Big{)}-\epsilon C_{F}x_{1}\Big{(}B(s_{2})+B_{c}(s_{2})\Big{)}\Big{]},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}B_{3}=$ $\frac{2x_{b}}{s_{2}(x+i\epsilon)(x_{1}+i\epsilon)}\frac{1}{\epsilon}\Big{[}C_{A}(-2+\epsilon)\Big{(}x_{2}B(s_{0})-xB(s_{2})\Big{)}+C_{F}(8+\epsilon^{2})x_{1}\Big{(}B(s_{2})+B_{c}(s_{2})\Big{)}\Big{]}.$ (69)
For $h_{4}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}A_{4}=$ $-\frac{2x_{2}(-2+\epsilon)}{s_{2}(x+i\epsilon)(x_{1}+i\epsilon)}\frac{1}{\epsilon}\Big{[}C_{A}x\Big{(}B(s_{0})-B(s_{2})\Big{)}-C_{F}\epsilon x_{1}\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}\Big{]},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}B_{4}=$ $\frac{2x_{b}}{s_{2}(x+i\epsilon)(x_{1}+i\epsilon)}\frac{1}{\epsilon}\Big{[}C_{A}(-2+\epsilon)\Big{(}x_{2}B(s_{0})-xB(s_{2})\Big{)}+C_{F}(8+\epsilon^{2})x_{1}\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}\Big{]}.$ (70)
For $h_{5}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}A_{5}=$ $-\frac{2x_{2}(2C_{A}+\epsilon C_{F})}{s_{2}(x_{1}+i\epsilon)}\Big{(}B(s_{0})-B(s_{2})\Big{)}$
$+\frac{2(C_{A}-2C_{F})x_{2}^{2}(-4+\epsilon^{2})}{s_{2}(x+i\epsilon)^{2}\epsilon}\Big{(}B(s_{1})-B(s_{2})\Big{)}$
$-\frac{2x_{2}}{s_{2}(x+i\epsilon)\epsilon}\Big{[}\epsilon(2C_{A}+\epsilon C_{F})B(s_{0})-2\Big{(}C_{A}\epsilon+C_{F}(4-3\epsilon+\epsilon^{2})\Big{)}B(s_{2})+C_{F}(8+\epsilon^{2})B_{c}(s_{2})\Big{]},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}B_{5}=$ $\frac{2x_{b}(2C_{A}+\epsilon C_{F})}{s_{2}(x_{1}+i\epsilon)}\Big{(}B(s_{0})-B(s_{2})\Big{)}$
$-\frac{2(C_{A}-2C_{F})x_{2}x_{b}}{s_{2}(x+i\epsilon)^{2}\epsilon}\Big{(}B(s_{1})-B(s_{2})\Big{)}$
$+\frac{2x_{b}}{s_{2}(x+i\epsilon)\epsilon}\Big{(}\epsilon(2C_{A}+\epsilon C_{F})B(s_{0})-2\epsilon(C_{A}-2C_{F})B(s_{2})+(8+\epsilon^{2})C_{F}B_{c}(s_{2})\Big{)}.$ (71)
For $h_{6}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}(A_{5}-A_{6})=$ $\frac{4C_{F}x_{2}(-2+\epsilon)}{s_{2}(x+i\epsilon)}B_{c}(s_{2}),$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}(B_{5}-B_{6})=$ $\frac{4C_{F}x_{b}(8+\epsilon^{2})}{s_{2}(x+i\epsilon)\epsilon}B_{c}(s_{2}).$ (72)
By using the formula
$\frac{1}{x+i\epsilon}=P\frac{1}{x}-i\pi\delta(x),$ (73)
$P\frac{1}{x}$ or $\delta(x)$ may contribute to the real parts of $A_{i},B_{i}$, since $B(s_{i})$ is complex for general $s_{i}$. The contribution proportional to $\delta(x)$ (or $P\frac{1}{x}$) is called pole contribution (or non-pole contribution). Pole contribution may be proportional to $\delta(x)$ or $\delta(x_{1})$. The former is called soft-gluon-pole(SGP) contribution, and the latter is called soft-fermion-pole(SFP) contribution. Before proceeding, we should show that SFP does not contribute to $h_{i}$, otherwise, the PV in eq.(25) for $x_{1}$ is ill-defined.
From the explicit results of $h_{i}$, $h_{3,4,5,6}$ may contain SFP contributions, which are given by following combination
$\frac{B(s_{0})-B(s_{2})}{x_{1}+i\epsilon},\ \text{or}\ \frac{x_{2}B(s_{0})-xB(s_{2})}{x_{1}+i\epsilon}.$ (74)
Under the limit $x_{1}\rightarrow 0$, we have $x=x_{2}$, $s_{0}=s_{2}$, so,
$\delta(x_{1})(B(s_{0})-B(s_{2}))=0,\ \delta(x_{1})(x_{2}B(s_{0})-xB(s_{2}))=0.$ (75)
Thus, SFP contribution vanishes, and all $h_{i}$’s are well defined at $x_{1}=0$ and PV in eq.(25) for $x_{1}$ is well defined.
Similarly, $g_{i}^{\mu\nu}$ is decomposed as
$g_{i}^{\mu\nu}=2g_{s}^{2}\text{Re}\Big{[}C_{i}t_{a}^{\mu\nu}+D_{i}t_{b}^{\mu\nu}\Big{]}\delta^{n}(k_{2}+k_{b}-q),$ (76)
with $i=0,1,\cdots,4$. Using eq.(30), we get the coefficients $C_{i},D_{i}$ as follows.
For $g_{0}$, we have
$C_{0}=D_{0}=0.$ (77)
For $g_{1}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C_{1}=$ $0,$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}D_{1}=$ $\frac{4C_{F}(2x_{2}-x)x_{b}}{s_{2}x_{1}}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}\frac{1}{x+i\epsilon}\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}.$ (78)
Because $B(s_{2})-B_{c}(s_{2})$ is purely imaginary, only SGP gives nonzero contribution. Nonpole contribution is zero. $1/x_{1}$ is a PV, which is introduced by EOM relations. For $g_{2}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C_{2}=$ $0,$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}D_{2}=$ $\frac{4C_{F}x_{b}}{s_{2}x_{1}}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}.$ (79)
$D_{2}$ has no pole contribution. Because $B(s_{2})-B_{c}(s_{2})$ is purely imaginary, we have $\text{Re}D_{2}=0$. As a result, $g_{2}=0$.
For $g_{3}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C_{3}=$ $\frac{2x_{2}}{s_{2}}\Big{[}-\frac{2C_{A}+\epsilon C_{F}}{x_{1}}\Big{(}B(s_{0})-B(s_{2})\Big{)}+\frac{(C_{A}-2C_{F})x_{2}}{(x+i\epsilon)^{2}}\frac{\epsilon^{2}-4}{\epsilon}\Big{(}B(s_{1})-B(s_{2})\Big{)}$
$+\frac{1}{x+i\epsilon}\Big{(}-(2C_{A}+\epsilon C_{F})B(s_{0})+2(C_{A}+C_{F}\frac{4-3\epsilon+\epsilon^{2}}{\epsilon})B(s_{2})+2C_{F}\frac{-4+\epsilon-\epsilon^{2}}{\epsilon}B_{c}(s_{2})\Big{)}\Big{]},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}D_{3}=$ $\frac{2x_{b}}{s_{2}}\Big{[}\frac{2C_{A}+\epsilon C_{F}}{x_{1}}\Big{(}B(s_{0})-B(s_{2})\Big{)}-\frac{(C_{A}-2C_{F})x_{2}}{(x+i\epsilon)^{2}}\frac{\epsilon^{2}-4}{\epsilon}\Big{(}B(s_{1})-B(s_{2})\Big{)}$
$+\frac{1}{x+i\epsilon}\Big{(}(2C_{A}+\epsilon C_{F})B(s_{0})+2(-C_{A}+C_{F}\frac{4+\epsilon+\epsilon^{2}}{\epsilon})B(s_{2})+2C_{F}\frac{-4+\epsilon-\epsilon^{2}}{\epsilon}B_{c}(s_{2})\Big{)}\Big{]}.$ (80)
For $g_{4}$, we notice that $g_{3}-g_{4}$ is very simple,
$\frac{1}{N_{c}C_{F}}p_{a}^{+}(C_{3}-C_{4})=$ $0,$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}(D_{3}-D_{4})=$ $\frac{8C_{F}x_{b}}{s_{2}(x+i\epsilon)}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}\Big{(}B(s_{2})-B_{c}(s_{2})\Big{)}.$ (81)
For convenience, we define $\Delta g^{\mu\nu}\equiv g_{3}^{\mu\nu}-g_{4}^{\mu\nu}$. Its expression is shown above. As can be seen, $\Delta g$ contains only SGP contribution.
As a summary, we find $g_{0}=g_{2}=0$. $g_{1}$ is nonzero, but contains only SGP contribution. Considering eq.(41), we have $\tilde{g}_{1}=0$. Thus, QCD gauge invariance is preserved. $g_{3},g_{4}$ contain SGP and non-pole contributions. Next, we present the results separately.
V.1.1 SGP contribution
The limit $x\rightarrow 0$ in $g_{i}$ gives SGP contribution. What is special in dimensional regularization is $B(s_{0})=0$ if $s_{0}=0$. In addition, $B(s_{1})$ should be expanded near $s_{2}$ since in physical region $s_{1}=s_{2}-s_{0}$.
$B(s_{1})-B(s_{2})=\frac{\epsilon}{2}\frac{s_{0}}{s_{2}}B(s_{2})+\frac{\epsilon}{4}(1+\frac{\epsilon}{2})\frac{s_{0}^{2}}{s_{2}^{2}}B(s_{2})+O(s_{0}^{3}).$ (82)
$s_{0}$ in right hand side eliminates the double pole $1/(x+i\epsilon)^{2}$ in $C_{3},D_{4}$. After this we get SGP contribution by the replacement $1/(x+i\epsilon)\rightarrow-i\pi\delta(x)$. The results are
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C_{1}^{SGP}=$ $0,$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}D_{1}^{SGP}=$ $\delta(x)\frac{16\pi C_{F}x_{b}}{s_{2}}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}\text{Im}\Big{(}B(s_{2})\Big{)},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C_{3}^{SGP}=$ $\delta(x)\frac{8\pi C_{F}x_{2}}{s_{2}}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}\text{Im}\Big{(}B(s_{2})\Big{)},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}D_{3}^{SGP}=$ $\delta(x)\frac{8\pi C_{F}x_{b}}{s_{2}}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}\text{Im}\Big{(}B(s_{2})\Big{)}.$ (83)
So,
$C_{3}^{SGP}/x_{2}=D_{3}^{SGP}/x_{b},\ D_{1}^{SGP}=2D_{3}^{SGP}.$ (84)
These two relations are important.
As discussed in Sec.III, it is $\frac{1}{2}g_{1}+g_{3}$ that gives the final contribution from $T_{F}(x_{2},x_{2})$. From above result, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}(-\frac{1}{2}C_{1}^{SGP}+C_{3}^{SGP})=$ $\frac{1}{N_{c}C_{F}}p_{a}^{+}C_{3}^{SGP},$
$\frac{1}{N_{c}C_{F}}p_{a}^{+}(-\frac{1}{2}D_{1}^{SGP}+D_{3}^{SGP})=$ $0.$ (85)
The vanishing of the second equation indicates SGP part has the same tensor structure as tree level.
For $g_{4}$, because $T_{\Delta}(x_{1},x_{2})=-T_{\Delta}(x_{2},x_{1})$, the SGP contribution does not exist.
In summary, SGP contribution to the nonderivative part of $W^{\mu\nu}$ is
$8N_{c}^{2}C_{F}W^{\mu\nu}_{SGP}=$ $-\frac{2g_{s}^{2}}{\pi}t_{a}^{\mu\nu}\int dk_{b}^{-}\int dk^{+}dk_{1}^{+}\bar{q}(x_{b})\delta^{n}(k+k_{1}+k_{b}-q)C_{3}^{SGP}T_{F}(x_{2},x_{2}).$ (86)
V.1.2 Non-pole contribution
For non-pole contribution, $x\neq 0$, and only the real part of $B(s_{i})$ contributes to $g_{i}$. The combination $B(s_{2})-B_{c}(s_{2})$ is purely imaginary and does not contribute. So, $g_{0,1,2}$ are zero. Moreover, the difference between $D_{3}$ and $D_{4}$ is proportional to $B(s_{2})-B_{c}(s_{2})$, thus the non-pole part of $g_{3}$ and $g_{4}$ is the same, i.e.,
$C_{4}^{NP}=C_{3}^{NP},\ D_{4}^{NP}=D_{3}^{NP}.$ (87)
For $C_{3}^{NP}$, we have
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C^{NP}_{3}=$ $-\frac{4x_{2}C_{A}}{x_{1}s_{2}}\text{Re}\Big{(}B(s_{0})-B(s_{2})\Big{)}-\frac{4x_{2}}{xs_{2}}\text{Re}\Big{(}C_{A}B(s_{0})-(C_{A}-2C_{F})B(s_{2})\Big{)}$
$-\frac{8(C_{A}-2C_{F})}{\epsilon}\frac{x_{2}^{2}}{s_{2}x^{2}}\text{Re}\Big{(}B(s_{1})-B(s_{2})\Big{)}.$ (88)
Because $s_{0},s_{1}$ can be negative or positive, $C_{3}^{NP}$ is non-zero. Note that
$\frac{1}{N_{c}C_{F}}p_{a}^{+}\Big{[}\frac{1}{x_{2}}C_{3}+\frac{1}{x_{b}}D_{3}\Big{]}=$ $\frac{8C_{F}}{s_{2}(x+i\epsilon)}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}(B(s_{2})-B_{c}(s_{2})),$ (89)
which does not contain non-pole contribution. So, the non-pole parts of $C_{3}$ and $D_{3}$ satisfy
$\frac{1}{x_{2}}C_{3}^{NP}+\frac{1}{x_{b}}D_{3}^{NP}=0.$ (90)
In summary, for non-derivative part of $W^{\mu\nu}$, the non-pole contribution is
$8N_{c}^{2}C_{F}W^{\mu\nu}_{NP}=$ $-\frac{2g_{s}^{2}}{\pi}\int dk_{b}^{-}\int dk^{+}dk_{1}^{+}\bar{q}(x_{b})\delta^{n}(k_{b}+k+k_{1}-q)\Big{(}C_{3}^{NP}t_{a}^{\mu\nu}+D_{3}^{NP}t_{b}^{\mu\nu}\Big{)}\Big{[}T_{F}(x_{1},x_{2})+T_{\Delta}(x_{1},x_{2})\Big{]}.$ (91)
Both $t_{a}^{\mu\nu}$ and $t_{b}^{\mu\nu}$ appear in this part, while only $t_{a}^{\mu\nu}$ appears in SGP part and in tree level result. Moreover, the coefficients here are divergent. Different from twist-2 cases, the new tensor structure with divergent coefficients does not imply a breaking of factorization, as we will explain later.
V.2 Corrections to derivative part of $W^{\mu\nu}$
Similar to tree level contribution, for this part, only $H_{4},H_{6}$ are nonzero and only SGP contribution is possible. We have
$8N_{c}^{2}C_{F}W_{de}^{\mu\nu}=$ $\int dk_{b}^{-}\int dk^{+}dk_{1}^{+}\delta(q^{-}-k_{b}^{-})\delta(k^{+}+k_{1}^{+}-q^{+})(H_{6}^{\mu\nu}-\frac{1}{2}H_{4}^{\mu\nu})M^{(1)}_{\gamma^{+},\partial G^{+}}$, (92)
and
$H_{6}^{\mu\nu}=H_{4}^{\mu\nu}=-i\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}].$ (93)
Because $k_{\perp}=k_{1\perp}=0$ in the trace, the trace is an on-shell quantity. After calculation, we have
$8N_{c}^{2}C_{F}W^{\mu\nu}_{de}=$ $-\frac{2g_{s}^{2}}{\pi}\int dk_{b}^{-}\int dk^{+}dk_{1}^{+}\delta(q^{-}-k_{b}^{-})\delta(k^{+}+k_{1}^{+}-q^{+})T_{F}(x_{2},x_{2})Eg_{\perp}^{\mu\nu}\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}.$ (94)
Both $\mu,\nu$ are transverse. The coefficient $E$ is related to $C_{3}^{SGP}$ as follows
$g_{s}^{2}E=g_{s}^{2}p_{a}\cdot qC_{3}^{SGP}=$ $\alpha_{s}\frac{N_{c}C_{F}^{2}}{x_{2}p_{a}^{+}}\delta(x)\Big{(}\frac{4\pi\mu^{2}}{Q^{2}}\Big{)}^{\epsilon/2}\frac{1}{\Gamma(1-\frac{\epsilon}{2})}\Big{[}\frac{8}{\epsilon^{2}}+\frac{6}{\epsilon}+8-\pi^{2}+O(\epsilon)\Big{]},$ (95)
which is just the correction to quark form factor, as pointed out in Chen et al. (2017). Note that $s_{2}=Q^{2}$ for virtual correction.
V.3 Total virtual corrections
Now, the complete virtual correction to hadronic tensor is the sum of eqs.(86,91,94), that is,
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $8N_{c}^{2}C_{F}[W^{\mu\nu}_{SGP}+W^{\mu\nu}_{NP}+W^{\mu\nu}_{de}]$
$=$ $-\frac{2g_{s}^{2}}{\pi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}\delta(k_{b}^{-}-q^{-})\delta(k_{2}^{+}-q^{+})\Big{\{}$
$\int dk^{+}T_{F}(x_{2},x_{2})\Big{[}\delta^{n-2}(q_{\perp})t_{a}^{\mu\nu}+\tilde{s}_{\perp}^{\rho}\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}g_{\perp}^{\mu\nu}p_{a}\cdot q\Big{]}C_{3}^{SGP}$
$+\int dk^{+}\delta^{n-2}(q_{\perp})\Big{(}C_{3}^{NP}t_{a}^{\mu\nu}+D_{3}^{NP}t_{b}^{\mu\nu}\Big{)}\Big{(}T_{F}(x_{1},x_{2})+T_{\Delta}(x_{1},x_{2})\Big{)}\Big{\}}.$ (96)
If there is no non-pole contribution, the virtual correction has the same structure as tree level hadronic tensor. The pole contribution, the last second line of above result, satisfies QED gauge invariance, as shown for tree level hadronic tensor. The correction to derivative part is the same as quark form factor, in agreement with the conclusion of Chen et al. (2017). In Chen et al. (2017), such correction is inferred from Ward identity for longitudinal gluon $G^{+}$. Here we recover the result by direct calculation. The non-pole contribution, the last line, also satisfies such invariance: with $q_{\perp}=0$,
$q_{\mu}\Big{(}C_{3,NP}t_{a}^{\mu\nu}+D_{3,NP}t_{b}^{\mu\nu}\Big{)}=$ $\tilde{s}_{\perp}^{\nu}p_{a}\cdot p_{b}\Big{(}x_{b}C_{3,NP}+x_{2}D_{3,NP}\Big{)}.$ (97)
Due to eq.(90), this is zero. Thus, QED gauge invariance is satisfied.
From $W^{\mu\nu}$, $I\langle L\rangle$ and $I\langle P_{7}\rangle$ can be obtained easily,
$I\langle L\rangle\Big{|}_{v}=$ $I\langle L\rangle\Big{|}_{v}^{SGP}+I\langle L\rangle\Big{|}_{v}^{NP},$
$I\langle P_{7}\rangle\Big{|}_{v}=$ $I\langle P_{7}\rangle\Big{|}_{v}^{SGP}.$ (98)
For SGP part, we have
$I\langle L\rangle\Big{|}^{SGP}_{v}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int\frac{d\xi}{\xi}\frac{dx_{b}}{x_{b}}\frac{dx_{2}}{x_{2}}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$
$\times(-32)(N_{c}^{2}-1)\Big{(}\frac{8}{\epsilon^{2}}+\frac{6}{\epsilon}+8-\pi^{2}\Big{)},$
$I\langle P_{7}\rangle\Big{|}^{SGP}_{v}=$ $-\frac{1}{4l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta}I\langle L\rangle\Big{|}^{SGP}_{v}.$ (99)
The overall factor $2$ in the first equation comes from the contribution of conjugated diagrams.
Only $I\langle L\rangle$ receives non-pole contribution:
$I\langle L\rangle\Big{|}^{NP}_{v}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int\frac{d\xi}{\xi}\frac{dx_{b}}{x_{b}}\frac{dx_{2}}{x_{2}}\bar{q}(x_{b})T_{F}(x^{*}_{2}-x_{2},x^{*}_{2})\delta(1-\hat{x}_{b})$
$\times\Big{(}128\hat{x}_{2}\Big{[}(\frac{1}{\epsilon}+1)\theta(x_{2}-x^{*}_{2})-\theta(x^{*}_{2}-x_{2})\frac{1}{2}\ln\frac{\hat{x}_{2}}{\hat{x}_{2}-1}\Big{]}$
$-\theta(-x_{2})\frac{64N_{c}^{2}}{1-\hat{x}_{2}}+64(1-N_{c}^{2}\theta(x_{2}))\Big{)}.$ (100)
As stated before, non-pole part is divergent. In these expressions,
$\bar{A}_{\epsilon}=\frac{1}{\Gamma(1-\frac{\epsilon}{2})}\Big{(}\frac{4\pi\mu^{2}}{Q^{2}}\Big{)}^{\epsilon/2},$ (101)
and
$\hat{x}_{2}\equiv\frac{x_{2}^{*}}{x_{2}},\ \hat{x}_{b}\equiv\frac{x_{b}^{*}}{x_{b}},\ x_{2}^{*}=\xi,\ x_{b}^{*}=\tau_{\xi},\ \xi=\frac{q^{+}}{p_{a}^{+}}.$ (102)
For SGP contribution, $x_{2}^{*}$ and $x_{b}^{*}$ serve as the lower bounds of integration about $x_{2}$, $x_{b}$, respectively. For non-pole contribution, the bounds for $x_{2}$ are determined by the theta functions and the support of $T_{F}$.
VI Real corrections
For real corrections, $q_{\perp}\neq 0$ is assumed. $W^{\mu\nu}$ thus depends on $p_{a}^{\mu},p_{b}^{\mu},q_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$. Because of QED gauge invariance $q_{\mu}W^{\mu\nu}=0$, there are six independent tensor structures for transversely polarized caseArnold et al. (2009). For our convenience, the tensors can be chosen as
$P_{1,\ldots,4}^{\mu\nu}=$ $\tilde{s}_{\perp}\cdot q_{\perp}\{g_{\perp}^{\mu\nu}-\frac{q_{\perp}^{\mu}q_{\perp}^{\nu}}{q_{\perp}^{2}},\ \tilde{p}_{a}^{\mu}\tilde{p}_{a}^{\nu},\tilde{p}_{b}^{\mu}\tilde{p}_{b}^{\nu},\tilde{p}_{a}^{\mu}\tilde{p}_{b}^{\nu}+\tilde{p}_{b}^{\mu}\tilde{p}_{a}^{\nu}\},$
$P_{5,6}^{\mu\nu}=$ $\{\tilde{p}_{a}^{\mu}\Big{(}\tilde{s}_{\perp}^{\nu}-\frac{\tilde{s}_{\perp}\cdot q_{\perp}}{q_{\perp}^{2}}q_{\perp}^{\nu}\Big{)}+(\mu\leftrightarrow\nu),\ \tilde{p}_{b}^{\mu}\Big{(}\tilde{s}_{\perp}^{\nu}-\frac{\tilde{s}_{\perp}\cdot q_{\perp}}{q_{\perp}^{2}}q_{\perp}^{\nu}\Big{)}+(\mu\leftrightarrow\nu)\},$ (103)
with $\tilde{p}_{i}^{\mu}=p_{i}^{\mu}-p_{i}\cdot qq^{\mu}/q^{2}$, so that $q\cdot\tilde{p}_{i}=0$.
$W^{\mu\nu}$ is decomposed as
$W^{\mu\nu}=\sum_{i=1}^{6}W_{i}P_{i}^{\mu\nu},$ (104)
with $W_{i}$ a scalar function of $p_{a},p_{b},q$. To solve $W_{i}$, we first contract both sides of eq.(104) with $P_{j}^{\mu\nu}$ to get
$P_{j}\cdot W=\sum_{i=1}^{6}W_{i}P_{i}\cdot P_{j}.$ (105)
The dot product represents the contraction of Lorentz indices, e.g., $P_{j}\cdot W=P_{j,\mu\nu}W^{\mu\nu}$. Eq.(105) can be solved directly. But the coefficients before $P_{i}\cdot W$ may depend on spin vector $\tilde{s}_{\perp}$ in a very complicated way, since $P_{i}\cdot P_{j}$ depends on $\tilde{s}_{\perp}$. Noticing that $W_{i}$ does not depend on the direction of $q_{\perp}$, we can integrate out the angles of $q_{\perp}$ on both sides of eq.(105) and then to solve the obtained equations. According to this treatment, we successfully get
$W_{i}=\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j},\ \xi_{j}\equiv\int d\Omega_{n-2}P_{j}\cdot W,$ (106)
where $d\Omega_{n-2}$ is for $q_{\perp}$. After integration, $\xi_{j}$ depends on $\tilde{s}_{\perp}^{2},q_{\perp}^{2},q\cdot p_{a},q\cdot p_{b}$. The coefficients $C_{ij}$ do not depend on $\tilde{s}_{\perp}$, whose expressions are
$C_{11}=$ $\frac{1}{(\epsilon-1)q_{t}^{2}},$
$C_{22}=$ $-\frac{\left(q\cdot p_{b}\right){}^{4}}{q_{t}^{6}\left(p_{a}\cdot p_{b}\right){}^{4}},$
$C_{23}=$ $-\frac{\left(q_{t}^{2}p_{a}\cdot p_{b}-q\cdot p_{a}q\cdot p_{b}\right){}^{2}}{q_{t}^{6}\left(p_{a}\cdot p_{b}\right){}^{4}},$
$C_{24}=$ $\frac{\left(q\cdot p_{b}\right){}^{2}\left(q_{t}^{2}p_{a}\cdot p_{b}-q\cdot p_{a}q\cdot p_{b}\right)}{q_{t}^{6}\left(p_{a}\cdot p_{b}\right){}^{4}},$
$C_{33}=$ $-\frac{\left(q\cdot p_{a}\right){}^{4}}{q_{t}^{6}\left(p_{a}\cdot p_{b}\right){}^{4}},$
$C_{34}=$ $\frac{\left(q\cdot p_{a}\right){}^{2}\left(q_{t}^{2}p_{a}\cdot p_{b}-q\cdot p_{a}q\cdot p_{b}\right)}{q_{t}^{6}\left(p_{a}\cdot p_{b}\right){}^{4}},$
$C_{44}=$ $-\frac{q_{t}^{4}\left(p_{a}\cdot p_{b}\right){}^{2}-2q_{t}^{2}p_{a}\cdot p_{b}q\cdot p_{a}q\cdot p_{b}+2\left(q\cdot p_{a}\right){}^{2}\left(q\cdot p_{b}\right){}^{2}}{2q_{t}^{6}\left(p_{a}\cdot p_{b}\right){}^{4}},$
$C_{55}=$ $\frac{\left(q\cdot p_{b}\right){}^{2}}{2(\epsilon-1)q_{t}^{2}\left(p_{a}\cdot p_{b}\right){}^{2}},$
$C_{56}=$ $\frac{q\cdot p_{a}q\cdot p_{b}-q_{t}^{2}p_{a}\cdot p_{b}}{2(\epsilon-1)q_{t}^{2}\left(p_{a}\cdot p_{b}\right){}^{2}},$
$C_{66}=$ $\frac{\left(q\cdot p_{a}\right){}^{2}}{2(\epsilon-1)q_{t}^{2}\left(p_{a}\cdot p_{b}\right){}^{2}}.$ (107)
Notice that $C_{ij}$ is symmetric, $C_{ij}=C_{ji}$. $q_{t}=|\vec{q}_{\perp}|$.
Then,
$I\langle L\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})L^{\mu\nu}W_{\mu\nu}$
$=$ $\sum_{i}\int d^{n}q\delta(q^{2}-Q^{2})L\cdot P_{i}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}.$ (108)
Since $C_{ij}$ and $\xi_{j}$ are functions of $q_{\perp}^{2}$, the angle of $q_{\perp}$ is just contained in $L\cdot P_{i}$. Again, we integrate out the angle of $q_{\perp}$ first,
$\int d^{n}q\delta(q^{2}-Q^{2})L\cdot P_{i}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}=$ $\int\frac{dq^{+}}{2q^{+}}\int dq_{t}q_{t}^{n-3}\Big{(}\int d\Omega_{n-2}L\cdot P_{i}\Big{)}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}.$ (109)
There is a transverse momentum $l_{\perp,cs}$ in $L^{\mu\nu}$, with lepton momentum $l^{\mu}$ given in eq.(8). Since $P_{i}$ contains only one $\tilde{s}_{\perp}$, we must have
$\Big{(}\int d\Omega_{n-2}L\cdot P_{i}\Big{)}\propto l_{\perp,cs}\cdot\tilde{s}_{\perp}\Omega_{n-2}.$ (110)
To obtain the coefficients is easy and we have
$\Big{(}\int d\Omega_{n-2}L\cdot P_{i}\Big{)}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}=l_{\perp,cs}\cdot\tilde{s}_{\perp}\sum_{i}a_{i}\xi_{i}\frac{1}{\tilde{s}_{\perp}^{2}}.$ (111)
Then,
$I\langle L\rangle=$ $\frac{l_{\perp,cs}\cdot\tilde{s}_{\perp}}{\tilde{s}_{\perp}^{2}}\sum_{i}\int\frac{dq^{+}}{2q^{+}}\int dq_{t}q_{t}^{n-3}a_{i}\xi_{i}.$ (112)
The coefficients $a_{i}$ are
$a_{1}=$ $0,$
$a_{2}=$ $-\frac{4QE_{t}^{2}l_{cs}^{z}}{q_{t}^{2}\left(q\cdot p_{a}\right){}^{2}},$
$a_{3}=$ $\frac{16Ql_{cs}^{z}\left(q\cdot p_{a}\right){}^{2}}{E_{t}^{2}q_{t}^{2}\left(p_{a}\cdot p_{b}\right){}^{2}},$
$a_{4}=$ $0,$
$a_{5}=$ $-\frac{4E_{t}l_{cs}^{z}}{q\cdot p_{a}},$
$a_{6}=$ $\frac{8l_{cs}^{z}q\cdot p_{a}}{E_{t}p_{a}\cdot p_{b}},$ (113)
with $E_{t}=\sqrt{Q^{2}-q_{\perp}^{2}}=\sqrt{Q^{2}+q_{t}^{2}}$, $l_{cs}^{z}=\frac{Q}{2}\cos\theta$. A nontrivial feature is all $a_{i}$ are proportional to $l^{z}_{cs}$ or $\cos\theta$.
The formula can be further simplified. With $\xi_{i}$ given in eq.(106), we have
$I\langle L\rangle=$ $\frac{l_{\perp,cs}\cdot\tilde{s}_{\perp}}{\tilde{s}_{\perp}^{2}}\sum_{i}\int\frac{dq^{+}}{2q^{+}}\int dq_{t}q_{t}^{n-3}a_{i}\int d\Omega_{n-2}P_{i}\cdot W$
$=$ $\frac{l_{\perp,cs}\cdot\tilde{s}_{\perp}}{\tilde{s}_{\perp}^{2}}\sum_{i}\int\frac{dq^{+}}{2q^{+}}\int dq_{t}q_{t}^{n-3}\int d\Omega_{n-2}a_{i}P_{i}\cdot W$
$=$ $\frac{l_{\perp,cs}\cdot\tilde{s}_{\perp}}{\tilde{s}_{\perp}^{2}}\sum_{i}\int d^{n}q\delta(q^{2}-Q^{2})a_{i}P_{i}\cdot W.$ (114)
Because
$P_{i}\cdot W=\tilde{s}_{\perp}^{\tau}\tilde{s}_{\perp\rho}P_{i,\mu\nu\tau}W^{\mu\nu,\rho},$ (115)
and
$\int d^{n}q\delta(q^{2}-Q^{2})a_{i}P_{i,\mu\nu\tau}W^{\mu\nu,\rho}=\tilde{W}_{i}g_{\perp\tau}^{\rho},\ \tilde{W}_{i}=\frac{1}{n-2}\int d^{n}q\delta(q^{2}-Q^{2})a_{i}P_{i,\mu\nu\tau}W^{\mu\nu,\rho}g_{\perp\rho}^{\tau},$ (116)
we have
$I\langle L\rangle=l_{\perp,cs}\cdot\tilde{s}_{\perp}\sum_{i=1}^{6}\tilde{W}_{i}.$ (117)
This is our main formula for real corrections. There is no $\tilde{s}_{\perp}$ or $l_{\perp,cs}$ in $\tilde{W}_{i}$ now. With $W^{\mu\nu}$ replaced by $\tilde{W}_{i}$, the formula eq.(40) can be applied. Further calculation is the same as that for real correction of weighted cross section studied in Chen et al. (2017).
Same as weighted cross section, there is no contribution from two-point correlation functions, i.e.,Fig.3. The amplitudes in these diagrams do not contain any absorptive part. So, $M^{(0)}_{\gamma_{\perp}\gamma_{5}}$ does not contribute and we just need to consider three-point distribution functions. Still, three types of poles contribute: SGP, SFP and hard pole(HP). The details can be found in Chen et al. (2017), here we just present the final result for each pole contribution. We note that for HP and SFP, the momentum fraction of initial gluon $x\neq 0$, and it can be shown easily that the collinear expansion based on $G^{+}$ or $G_{\perp}$ leads to the same hard coefficientsEguchi et al. (2007). We have checked this by direct calculation. In App.B, we present some details for the calculation of HP contribution based on $G_{\perp}$ expansion. The procedure based on $G^{+}$ expansion is too lengthy and not shown. For SGP contribution, we use $G^{+}$ to do collinear expansion only. Especially, for Fig.4, the contribution from $M^{(1)}_{\gamma_{\perp}\gamma_{5}}$ vanishes, because corresponding $W^{\mu\nu}$ projected by $\gamma_{5}\gamma_{\perp}$ is anti-symmetric in $\mu,\nu$. $M^{(1)}_{\gamma_{\perp}}$ does not contribute because
$M^{(1)}_{\gamma_{\perp}}(k^{+},k_{1}^{+})\Big{|}_{k^{+}=0}=0,$ (118)
which can be shown from PT symmetry or from the second equation of eq.(25) and eq.(31). In this work, we consider only the contribution proportional to $\bar{q}\otimes T_{F}$. Because $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}$ and $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$ are related to $T_{\Delta}$ rather than $T_{F}$, we do not consider the contribution from these two matrix elements in the following. Our SGP contributions are given by $M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}$ and $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$. As illustrated in Sec.III, SGP contributions from these two matrix elements can be expressed by $T_{F}$. In addition, we also use $L^{\mu\nu}$ instead of the projection operators $P_{i}^{\mu\nu}$ to calculate. The same $I\langle L\rangle$ is obtained. This is a check of our calculation.
Figure 3: The diagrams for the hard part of real corrections related to twist-3 two-point correlation functions. The conjugated diagram of (c) is not shown. The black dot represents quark-photon interaction.
VI.1 SGP contribution
SGP contribution is given by Fig.4. The short bar indicates the propagator is on-shell. Fig.4(c,d) are mirror diagramsEguchi et al. (2007), which do not contribute. Our calculation confirms this.
Figure 4: Diagrams for the hard part of SGP contributions in $\bar{q}+qg$ channel.(a-d) are of the left parts and the last two diagrams are of the right part. The black dot represents photon-quark interaction. The propagator with short bar is on-shell. Conjugated diagrams are not shown.
The results from Fig.4(a,b) are written as
$\{I\langle P_{7}\rangle,I\langle L\rangle\}=$ $2\times\frac{\alpha_{s}\{1,l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}$
$g_{0}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})+g_{1}\delta(1-\hat{x}_{b})+g_{2}\delta(1-\hat{x}_{2})+\frac{g_{3}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big{]},$ (119)
with
$\int_{r}\equiv\int_{\tau}^{1}\frac{d\xi}{\xi}\int_{\xi}^{1}\frac{dx_{2}}{x_{2}}\int_{\tau_{\xi}}^{1}\frac{dx_{b}}{x_{b}}.$ (120)
The integration bounds are determined by $k_{g}^{\pm}\geq 0$ with $k_{g}$ the momentum of final gluon.
For $I\langle P_{7}\rangle$, the result is
$g_{0}=$ $\frac{64}{\epsilon^{2}}+O\left(\epsilon^{1}\right),$
$g_{1}=$ $-\frac{16\left(\hat{x}_{2}^{2}+1\right)}{\left(1-\hat{x}_{2}\right)_{+}\epsilon}+8\left(\left(L_{2}-1\right)\hat{x}_{2}^{2}+L_{2}-\hat{x}_{2}\right)+O\left(\epsilon^{1}\right),$
$g_{2}=$ $-\frac{16\left(\hat{x}_{b}^{3}+\hat{x}_{b}\right)}{\epsilon\left(1-\hat{x}_{b}\right)_{+}}+8\left(L_{1}\hat{x}_{b}^{3}+\left(L_{1}+1\right)\hat{x}_{b}+1\right)+O\left(\epsilon^{1}\right),$
$g_{3}=$ $8\Big{(}\hat{x}_{2}^{2}\left(2\hat{z}-1\right)+\hat{x}_{2}\left(2\hat{z}^{2}-3\hat{z}+1\right)+\hat{z}\left(\hat{z}^{2}-2\hat{z}+2\right)\Big{)},$ (121)
with
$\hat{z}\equiv 1-\hat{x}_{2}(1-\hat{x}_{b})\text{ or }\hat{x}_{b}=\frac{\hat{x}_{2}+\hat{z}-1}{\hat{x}_{2}}.$ (122)
For $I\langle L\rangle$, the result is
$g_{0}=$ $-\frac{256}{\epsilon^{2}}+O\left(\epsilon^{1}\right),$
$g_{1}=$ $\frac{64\left(\hat{x}_{2}^{2}+1\right)}{\left(1-\hat{x}_{2}\right)_{+}\epsilon}-32\left(\left(L_{2}+1\right)\hat{x}_{2}^{2}+L_{2}-\hat{x}_{2}\right)+O\left(\epsilon^{1}\right),$
$g_{2}=$ $\frac{64\left(\hat{x}_{b}^{3}+\hat{x}_{b}\right)}{\epsilon\left(1-\hat{x}_{b}\right)_{+}}-32\left(L_{1}\hat{x}_{b}^{3}+\left(L_{1}+1\right)\hat{x}_{b}-1\right)+O\left(\epsilon^{1}\right),$
$g_{3}=$ $-\frac{32\left(\hat{x}_{2}+\hat{z}-1\right)\left(\hat{x}_{2}^{2}\left(\hat{z}\left(E_{t}-Q\right)+Q\right)+\hat{x}_{2}\left(\hat{z}-1\right)\hat{z}\left(E_{t}+Q\hat{z}\right)+Q\hat{x}_{2}^{3}\left(\hat{z}-1\right)+Q\hat{z}^{2}\left(\hat{z}^{2}-\hat{z}+1\right)\right)}{Q\hat{x}_{2}\hat{z}^{2}},$ (123)
with
$L_{1}=\Big{(}\frac{\ln(1-\hat{x}_{b})}{1-\hat{x}_{b}}\Big{)}_{+}-\frac{\ln\hat{x}_{b}}{1-\hat{x}_{b}},\ L_{2}=\Big{(}\frac{\ln(1-\hat{x}_{2})}{1-\hat{x}_{2}}\Big{)}_{+}.$ (124)
VI.2 HP contribution
The hard pole contribution is given by Figs.5,6. There are two channels which correspond to two different subprocesses: $\bar{q}+qg\rightarrow g+\gamma^{*}$ and $\bar{q}+q\bar{q}\rightarrow q+\gamma^{*}$. We next give their results separately.
Figure 5: Diagrams for the hard part of hard pole contributions from $\bar{q}+qg$ channel. (a,b,c) are of the left part and the last two diagrams are of the right part. Conjugated diagrams are not shown.
VI.2.1 $\bar{q}+qg$ channel
The diagrams are given in Fig.5. The pole condition is $(k_{1}-q)^{2}=0$ or
$x_{1}=x_{1}^{*}=\frac{Q^{2}}{2p_{a}\cdot q}.$ (125)
The results are summarized as
$\{I\langle P_{7}\rangle,I\langle L\rangle\}=$ $2\times\frac{\alpha_{s}\{1,l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\Big{[}$
$g_{0}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})+g_{1}\delta(1-\hat{x}_{b})+g_{2}\delta(1-\hat{x}_{2})+\frac{g_{3}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big{]}.$ (126)
For $I\langle P_{7}\rangle$, the result is
$g_{0}=$ $-\frac{64N_{c}{}^{2}}{\epsilon^{2}}-\frac{32N_{c}{}^{2}}{\epsilon}-16N_{c}{}^{2}+O\left(\epsilon^{1}\right),$
$g_{1}=$ $\frac{16\left(\left(\hat{x}_{2}+1\right)N_{c}{}^{2}\right)}{\left(1-\hat{x}_{2}\right)_{+}\epsilon}+\frac{8\left(\left(\hat{x}_{2}+1\right)\left(L_{2}\hat{x}_{2}-L_{2}+1\right)N_{c}{}^{2}\right)}{(1-\hat{x}_{2})_{+}}+O\left(\epsilon^{1}\right),$
$g_{2}=$ $\frac{16\left(\left(\hat{x}_{b}^{2}+1\right)\left(\hat{x}_{b}+N_{c}{}^{2}-1\right)\right)}{\epsilon\left(1-\hat{x}_{b}\right)_{+}}+\frac{8\left(\left(L_{1}\hat{x}_{b}^{3}-L_{1}\hat{x}_{b}^{2}+\left(L_{1}+2\right)\hat{x}_{b}-L_{1}\right)\left(\hat{x}_{b}+N_{c}{}^{2}-1\right)\right)}{(1-\hat{x}_{b})_{+}}+O\left(\epsilon^{1}\right),$
$g_{3}=$ $-8\left(\left(\hat{x}_{2}+\hat{z}^{2}\right)\left(N_{c}{}^{2}+\hat{z}-1\right)\right)+O\left(\epsilon^{1}\right).$ (127)
For $I\langle L\rangle$, the result is
$g_{0}=$ $\frac{256N_{c}{}^{2}}{\epsilon^{2}}+O\left(\epsilon^{1}\right),$
$g_{1}=$ $-\frac{64\left(\hat{x}_{2}+1\right)N_{c}{}^{2}}{\left(1-\hat{x}_{2}\right)_{+}\epsilon}+32\left(L_{2}\hat{x}_{2}+L_{2}-1\right)N_{c}{}^{2}+O\left(\epsilon^{1}\right),$
$g_{2}=$ $-\frac{64\left(\hat{x}_{b}^{2}+1\right)\left(\hat{x}_{b}+N_{c}{}^{2}-1\right)}{\epsilon\left(1-\hat{x}_{b}\right)_{+}}+32\left(L_{1}\hat{x}_{b}^{2}+L_{1}+2\right)\left(\hat{x}_{b}+N_{c}{}^{2}-1\right)+O\left(\epsilon^{1}\right),$
$g_{3}=$ $\frac{32\left(\hat{x}_{2}\hat{z}E_{t}+\left(\hat{z}-1\right)\hat{z}E_{t}+Q\hat{x}_{2}^{2}\right)\left(N_{c}{}^{2}+\hat{z}-1\right)}{\hat{x}_{2}E_{t}}+O\left(\epsilon^{1}\right).$ (128)
Except for $g_{0}$, the divergent part of $I\langle L\rangle$ is “4-4” times of the divergent part of $I\langle P_{7}\rangle$.
VI.2.2 $\bar{q}+q\bar{q}$ channel
This type of hard pole contribution is given by Fig.6.
Figure 6: Diagrams for the hard part of hard pole contributions from $\bar{q}+q\bar{q}$ channel. (a,b,c) are of the left part and the last two diagrams are of the right part. Conjugated diagrams are not shown.
The result is written as
$\{I\langle P_{7}\rangle,I\langle L\rangle\}=$ $2\times\frac{\alpha_{s}\{1,l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})$
$\Big{\{}T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})\Big{[}g_{0}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})+g_{1}\delta(1-\hat{x}_{b})+g_{2}\delta(1-\hat{x}_{2})+\frac{g_{3}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big{]}$
$+T_{F}(x_{2}-x_{1}^{*},-x_{1}^{*})\frac{g_{4}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big{\}}.$ (129)
Fig.6(a,c) gives $T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})$, while Fig.6(b,d) gives $T_{F}(x_{2}-x_{1}^{*},-x_{1}^{*})$. For the latter, there is no contribution proportional to delta functions.
For $I\langle P_{7}\rangle$, the result is
$g_{0}=$ $0,$
$g_{1}=$ $-\frac{32\hat{x}_{2}-16}{\epsilon}+8\left(\left(2\hat{x}_{2}-1\right)\left(\log\left(1-\hat{x}_{2}\right)-1\right)\right),$
$g_{2}=$ $0,$
$g_{3}=$ $-\frac{8(1-\hat{x}_{2})}{\hat{z}}\Big{[}-N_{c}\left(\hat{z}-1\right)\left(-2\hat{x}_{2}\left(\hat{z}-2\right)+\hat{z}^{2}-4\hat{z}+2\right)+\hat{z}\left(-2\hat{x}_{2}+2\hat{z}-1\right)\Big{]},$
$g_{4}=$ $-\frac{8(1-\hat{x}_{2})(1-\hat{x}_{b})}{\hat{z}}\Big{[}\left(2\hat{x}_{2}\left(\hat{z}-2\right)-\hat{z}^{2}+4\hat{z}-2\right)N_{c}+\left(\hat{z}-1\right)\hat{z}\left(2\hat{x}_{2}-\hat{z}+1\right)\Big{]}.$ (130)
For $I\langle L\rangle$, the result is
$g_{0}=$ $0,$
$g_{1}=$ $-\frac{64}{\epsilon}-\left(32-32\log\left(1-\hat{x}_{2}\right)\right),$
$g_{2}=$ $0,$
$g_{3}=$ $-\frac{32Q^{3}\left(\hat{x}_{2}-1\right)\hat{x}_{2}\left(\hat{z}-1\right)\left(2\hat{z}N_{c}-2N_{c}+\hat{z}\right)}{E_{t}^{3}\left(\hat{x}_{2}+\hat{z}-1\right)}+\frac{32\left(\hat{x}_{2}-1\right)\left(\hat{x}_{2}+\hat{z}-1\right)\left(\hat{z}^{2}N_{c}-\hat{z}\left(3N_{c}+1\right)+2N_{c}\right)}{\hat{x}_{2}\hat{z}}$,
$g_{4}=$ $-\frac{32\left(\hat{x}_{2}-1\right)\left(\hat{z}-1\right)\left(\hat{x}_{2}+\hat{z}-1\right)\left(\hat{z}\left(N_{c}-1\right)-2N_{c}+\hat{z}^{2}\right)}{\hat{x}_{2}\hat{z}}+\frac{32Q^{3}\left(\hat{x}_{2}-1\right)\hat{x}_{2}\left(\hat{z}-1\right)^{2}\left(2N_{c}+\hat{z}\right)}{E_{t}^{3}\left(\hat{x}_{2}+\hat{z}-1\right)}.$ (131)
It is noted that the divergent part of $g_{1}$ is very different for the two observables.
VI.3 SFP contributions
In this work, we concentrate on the contribution of $\bar{q}\otimes T_{F}$. For this case, the diagrams giving SFP are shown in Fig.7. Mirror diagrams are not shown, which do not contributeEguchi et al. (2007).
Figure 7: Diagrams for the hard part of SFP contribution from $\bar{q}+qg$ channel. (a,b) are of the left part and the last two diagrams are of the right part. Conjugated diagrams are not shown.
The result is finite.
$\{I\langle P_{7}\rangle,I\langle L\rangle\}=$ $2\times\frac{\alpha_{s}\{1,l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(0,x_{2})\frac{g_{3}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}.$ (132)
For $I\langle P_{7}\rangle$, the result is
$g_{3}=$ $-8\left(\left(\hat{x}_{2}-1\right)\left(\hat{z}-1\right)\left(-4\hat{x}_{2}\hat{z}+2\hat{x}_{2}^{2}+\hat{x}_{2}+\hat{z}^{2}\right)\right);$ (133)
For $I\langle L\rangle$, the result is
$g_{3}=$ $-\frac{32\left(\left(\hat{x}_{2}-1\right)\left(\hat{z}-1\right)\left(\hat{x}_{2}+\hat{z}-1\right)\left(\hat{x}_{2}E_{t}+2Q\hat{x}_{2}^{2}-Q\hat{z}^{2}\right)\right)}{Q\hat{x}_{2}\hat{z}}$. (134)
So far, we have listed all unsubtracted hard coefficients relevant to $\bar{q}\otimes T_{F}$. For $I\langle P_{7}\rangle$, all results are the same as our previous resultChen et al. (2017).
VII Renormalization and Subtraction
The counter terms and renormalization constants are given in Appendix.A. Here we present some details for the renormalization of twist-3 distribution functions. It is noted that in the definitions of $T_{F}$ and $M^{(1)}$ there is one explicit $g_{s}$. We should be careful about the renormalization of this coupling. Consider the renormalization for $W^{\mu\nu}$. First, we use bare quantities to write it as
$W^{\mu\nu}=\Gamma_{B}\otimes\bar{M}^{(1)}_{B}\otimes\bar{q}_{B},$ (135)
where $\bar{M}^{(1)}_{B}$ does not include $g_{s}^{B}$ explicitly. Then, the renormalization for $\Gamma_{B}$ is
$\Gamma_{B}=Z_{2}^{-4/2}Z_{3}^{-1/2}\Gamma_{R}.$ (136)
Now, we extract one $g_{s}^{B}$ from $\Gamma_{B}$ to define $\tilde{\Gamma}_{B}$ as follows
$\Gamma_{B}=\tilde{\Gamma}_{B}g_{s}^{B}.$ (137)
Renormalized $\Gamma_{R}$ and $\tilde{\Gamma}_{R}$ are defined similarly,
$\Gamma_{R}=\tilde{\Gamma}_{R}g_{s}.$ (138)
Then,
$\tilde{\Gamma}_{B}=Z_{2}^{-4/2}Z_{3}^{-1/2}\frac{g_{s}}{g_{s}^{B}}\tilde{\Gamma}_{R}.$ (139)
Using
$g_{s}^{B}=Z_{g}g_{s},\ Z_{g}=Z_{1F}Z_{2}^{-1}Z_{3}^{-1/2},$ (140)
we get
$\tilde{\Gamma}_{B}=Z_{2}^{-1}Z_{1F}^{-1}\tilde{\Gamma}_{R}.$ (141)
It is for $g_{s}^{B}\bar{M}_{B}^{(1)}$ that is renormalized as a whole,
$M_{B}^{(1)}=g_{s}^{B}\bar{M}^{(1)}_{B}=Z^{\prime}_{pdf}\otimes g_{s}\bar{M}_{R}^{(1)}=Z^{\prime}_{pdf}\otimes M_{R}^{(1)}$ (142)
$Z^{\prime}_{pdf}$ is the renormalizaition constant for twist-3 PDF, which is related to the evolution kernel. For twist-2 PDF,
$\bar{q}_{B}=Z_{pdf}\otimes\bar{q}_{R}.$ (143)
So,
$W^{\mu\nu}=$ $\tilde{\Gamma}_{B}\otimes g_{s}^{B}\bar{M}^{(1)}_{B}\otimes\bar{q}_{B}$
$=$ $\tilde{\Gamma}_{R}Z_{2}^{-1}Z_{1F}^{-1}\otimes[Z^{\prime}_{pdf}\otimes M^{(1)}_{R}]\otimes[Z_{pdf}\otimes\bar{q}_{R}].$ (144)
$\tilde{\Gamma}_{R}$ is calculated by using counter terms. At one loop level, it is
$\tilde{\Gamma}_{R}=\tilde{\Gamma}_{vir}+\tilde{\Gamma}_{tree}(1+2\delta z_{1}^{\gamma}+\delta z_{1F}-\delta z_{2}).$ (145)
We ignore real corrections to $\tilde{\Gamma}$ here. The last term $\delta z_{2}$ in $(\cdots)$ is from the counter term contribution to Fig.2(b). Considering eq.(144), we have
$W^{\mu\nu}=$ $[\tilde{\Gamma}_{vir}+\tilde{\Gamma}_{tree}(1+2\delta z_{1}^{\gamma}-2\delta z_{2})]\otimes[Z^{\prime}_{pdf}\otimes M^{(1)}_{R}]\otimes[Z_{pdf}\otimes\bar{q}_{R}].$ (146)
Note that $\delta z_{1}^{\gamma}=\delta z_{2}$. So,
$W^{\mu\nu}=$ $[\tilde{\Gamma}_{vir}+\tilde{\Gamma}_{tree}]\otimes[Z^{\prime}_{pdf}\otimes M^{(1)}_{R}]\otimes[Z_{pdf}\otimes\bar{q}_{R}].$ (147)
That is, wave function renormalization and counter term contributions cancel each other. As a net effect, the hard part is not affected by counter terms and self-energy corrections to external legs. Only distribution functions need renormalization. Divergence in hard part now is of IR type.
Next, we should make a collinear subtraction to remove the collinear divergence in the hard partCollins and Rogers (2008). The subtraction term is obtained by following replacement in tree level cross section,
$\bar{q}(x)\rightarrow\bar{q}(x)-\Delta\bar{q}(x),\ T_{F}(x,x)\rightarrow T_{F}(x,x)-\Delta T_{F}(x,x).$ (148)
$\Delta\bar{q}$ and $\Delta T_{F}$ are obtained from $Z_{pdf}$ and $Z^{\prime}_{pdf}$ in eq.(147). They are also related to the evolution kernels of $\bar{q}(x)$ and $T_{F}(x_{1},x_{2})$. In $\overline{\text{MS}}$ scheme, their expressions are
$\Delta\bar{q}(x)=$ $-\frac{\alpha_{s}}{2\pi}\Big{(}\frac{2}{\epsilon}-\gamma_{E}+\ln 4\pi\Big{)}\int_{x}^{1}\frac{dz}{z}P_{qq}(z)\bar{q}(\frac{x}{z}),$
$\Delta T_{F}(x,x)=$ $-\frac{\alpha_{s}}{2\pi}\Big{(}\frac{2}{\epsilon}-\gamma_{E}+\ln 4\pi\Big{)}\mathcal{F}_{q}\otimes T_{F}(x).$ (149)
$P_{qq}(z)$ is the standard DGLAP kernel,
$P_{qq}(z)=C_{F}\Big{[}\frac{1+z^{2}}{(1-z)_{+}}+\frac{3}{2}\delta(1-z)\Big{]}.$ (150)
The kernel for $T_{F}$ is a little complicatedKang and Qiu (2009); Braun et al. (2009); Zhou et al. (2010); Ma and Wang (2012); Schafer and Zhou (2012); Kang and Qiu (2012), which reads
$\mathcal{F}_{q}\otimes T_{F}(x)=$ $-N_{c}T_{F}(x,x)+\int_{x}^{1}\frac{dz}{z}\Big{[}P_{qq}(z)T_{F}(\xi,\xi)+\frac{N_{c}}{2}\Big{(}T_{\Delta}(x,\xi)+\frac{(1+z)T_{F}(x,\xi)-(1+z^{2})T_{F}(\xi,\xi)}{1-z}\Big{)}$
$+\frac{1}{2N_{c}}\Big{(}(1-2z)T_{F}(x,x-\xi)+T_{\Delta}(x,x-\xi)\Big{)}-\frac{1}{2}\frac{(1-z)^{2}+z^{2}}{\xi}T_{G+}(\xi,\xi)\Big{]},$ (151)
with $\xi=x/z$. $T_{G+}$ is a pure gluon twist-3 distribution function, which is ignored in the following calculation, since we are interested in quark contribution only.
From eq.(52), the tree level $I\langle P_{7}\rangle$ and $I\langle L\rangle$ can be obtained,
$I\langle P_{7}\rangle\Big{|}_{tree}=$ $-\frac{1}{2N_{c}}\int_{\tau}^{1}\frac{d\xi}{\xi}\bar{q}(x_{b}^{*})T_{F}(x_{2}^{*},x_{2}^{*}),$
$I\langle L\rangle\Big{|}_{tree}=$ $\tilde{s}_{\perp}\cdot l_{\perp,cs}\cos\theta\frac{2}{N_{c}}\int_{\tau}^{1}\frac{d\xi}{\xi}\bar{q}(x_{b}^{*})T_{F}(x_{2}^{*},x_{2}^{*}),$ (152)
with $x_{b}^{*}=\tau_{\xi}$ and $x_{2}^{*}=\xi$. The two tree level results differ by an $\epsilon$ independent constant factor. Their subtraction terms have the same relation. If the factorization is right, we must have
$I\langle w\rangle\Big{|}_{r+v}-I\langle w\rangle\Big{|}_{sub}=\text{finite.}$ (153)
for $w=P_{7},L$. We consider the subtraction for $I\langle P_{7}\rangle$ first.
VII.1 $I\langle P_{7}\rangle$ subtraction
The subtraction term is
$I\langle P_{7}\rangle\Big{|}_{sub}=$ $-\frac{1}{2N_{c}}\int\frac{d\xi}{\xi}\Big{[}(\Delta\bar{q}(x_{b}^{*}))T_{F}(x_{2}^{*},x_{2}^{*})+\bar{q}(x_{b}^{*})\Delta T_{F}(x_{2}^{*},x_{2}^{*})\Big{]}.$ (154)
With some transformations, it can be written into the standard form
$I\langle P_{7}\rangle\Big{|}_{sub}=$ $2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})\Big{\{}T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}+\delta(1-\hat{x}_{b})g_{1}^{(a)}(\hat{x}_{2})+\delta(1-\hat{x}_{2})g_{2}(\hat{x}_{b})\Big{]}$
$+T_{F}(x_{2}^{*},x_{2})\Big{[}\delta(1-\hat{x}_{b})g_{1}^{(b)}(\hat{x}_{2})\Big{]}+T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\Big{[}\delta(1-\hat{x}_{b})g_{1}^{(c)}(\hat{x}_{2})\Big{]}\Big{\}},$ (155)
with
$g_{0}=$ $\frac{-48+16N_{c}^{2}}{\epsilon}+(24-8N_{c}^{2})\ln\frac{\mu^{2}}{Q^{2}},$
$g_{1}^{(a)}(\hat{x}_{2})=$ $-\frac{16(1+\hat{x}_{2}^{2})}{\epsilon(1-\hat{x}_{2})_{+}}+\frac{8(1+\hat{x}_{2}^{2})}{(1-\hat{x}_{2})_{+}}\ln\frac{\mu^{2}}{Q^{2}},$
$g_{1}^{(b)}(\hat{x}_{2})=$ $\frac{16N_{c}^{2}(1+\hat{x}_{2})}{\epsilon(1-\hat{x}_{2})_{+}}-\frac{8N_{c}^{2}(1+\hat{x}_{2})}{(1-\hat{x}_{2})_{+}}\ln\frac{\mu^{2}}{Q^{2}},$
$g_{1}^{(c)}(\hat{x}_{2})=$ $\frac{16-32\hat{x}_{2}}{\epsilon}+(16\hat{x}_{2}-8)\ln\frac{\mu^{2}}{Q^{2}},$
$g_{2}(\hat{x}_{b})=$ $(N_{c}^{2}-1)\frac{16(1+\hat{x}_{b}^{2})}{\epsilon(1-\hat{x}_{b})_{+}}-(N_{c}^{2}-1)\frac{8(1+\hat{x}_{b}^{2})}{(1-\hat{x}_{b})_{+}}\ln\frac{\mu^{2}}{Q^{2}}.$ (156)
On the other hand, the complete virtual corrections is
$I\langle P_{7}\rangle\Big{|}_{v}=$ $2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big{]},$
$g_{0}=$ $8(N_{c}^{2}-1)\Big{(}\frac{8}{\epsilon^{2}}+\frac{6}{\epsilon}+8-\pi^{2}\Big{)}.$ (157)
Complete real corrections can be read from eqs.(121,127,130,133). Here we list their divergent parts for convenience.
$I\langle P_{7}\rangle\Big{|}_{r}\doteq$ $2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(a)}+\delta(1-\hat{x}_{b})g_{1}^{(a)}+\delta(1-\hat{x}_{2})g_{2}^{(a)}\Big{]}$
$+$ $\bar{q}(x_{b})T_{F}(x^{*}_{1},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(b)}+\delta(1-\hat{x}_{b})g_{1}^{(b)}+\delta(1-\hat{x}_{2})g_{2}^{(b)}\Big{]}$
$+$ $\bar{q}(x_{b})T_{F}(x^{*}_{1},x^{*}_{1}-x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(c)}+\delta(1-\hat{x}_{b})g_{1}^{(c)}+\delta(1-\hat{x}_{2})g_{2}^{(c)}\Big{]}\Big{\}},$ (158)
with
$g_{0}^{(a)}=$ $\frac{64}{\epsilon^{2}},$
$g_{1}^{(a)}=$ $-\frac{16(\hat{x}_{2}^{2}+1)}{\epsilon(1-\hat{x}_{2})_{+}},$
$g_{2}^{(a)}=$ $-\frac{16(\hat{x}_{b}+\hat{x}_{b}^{3})}{\epsilon(1-\hat{x}_{b})_{+}};$ (159)
and
$g_{0}^{(b)}=$ $\frac{-64N_{c}^{2}}{\epsilon^{2}}-\frac{32N_{c}^{2}}{\epsilon},$
$g_{1}^{(b)}=$ $\frac{16N_{c}^{2}(\hat{x}_{2}+1)}{\epsilon(1-\hat{x}_{2})_{+}},$
$g_{2}^{(b)}=$ $\frac{16(\hat{x}_{b}^{2}+1)(\hat{x}_{b}+N_{c}^{2}-1)}{\epsilon(1-\hat{x}_{b})_{+}};$ (160)
and
$g_{0}^{(c)}=$ $0,$
$g_{1}^{(c)}=$ $-\frac{32\hat{x}_{2}-16}{\epsilon},$
$g_{2}^{(c)}=$ $0.$ (161)
For $g_{0}$, the sum of virtual and real corrections is
$g_{0}|_{v+r}=g_{0}|_{v}+g_{0}|_{r}^{(a+b+c)}=(N_{c}^{2}-1)(\frac{64}{\epsilon^{2}}+\frac{48}{\epsilon})+\frac{64}{\epsilon^{2}}-\frac{64}{\epsilon^{2}}N_{c}^{2}-\frac{32}{\epsilon}N_{c}^{2}=\frac{16N_{c}^{2}-48}{\epsilon}.$ (162)
This is the same as $g_{0}|_{sub}$, given in eq.(155). For $I\langle P_{7}\rangle$, virtual correction contains vanishing $g_{1}$ and $g_{2}$, while the divergence of $g_{1},g_{2}$ from real correction is the same as that of the subtraction term, which can be seen from eq.(158) and eq.(155).
Now it is clear that
$I\langle P_{7}\rangle\Big{|}_{r}+I\langle P_{7}\rangle\Big{|}_{v}-I\langle P_{7}\rangle\Big{|}_{sub}=\text{finite.}$ (163)
So, $I\langle P_{7}\rangle$ can be factorized.
VII.2 $I\langle L\rangle$ subtraction
Since
$I\langle L\rangle|_{tree}=-4l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta I\langle P_{7}\rangle|_{tree},$ (164)
the subtraction terms have the same relation, i.e.,
$I\langle L\rangle|_{sub}=-4l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta I\langle P_{7}\rangle|_{sub}.$ (165)
However, virtual corrections do not have the same relation. We have
$I\langle L\rangle\Big{|}_{v}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta 2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{\tau}^{1}\frac{d\xi}{\xi}\int_{\tau_{\xi}}^{1}\frac{dx_{b}}{x_{b}}\int\frac{dx_{2}}{x_{2}}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big{]}$
$+\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\Big{[}g_{1}^{(c)}\theta(x_{2}-x_{2}^{*})+\text{finite.}\Big{]}\Big{\}},$ (166)
with
$g_{0}=$ $-32(N_{c}^{2}-1)\Big{(}\frac{8}{\epsilon^{2}}+\frac{6}{\epsilon}+8-\pi^{2}\Big{)},$
$g_{1}^{(c)}=$ $128\hat{x}_{2}(\frac{1}{\epsilon}+1).$ (167)
$g_{1}^{(c)}$ is given by nonpole contributions. Other finite terms can be obtained from eq.(99). The boundary for $x_{2}$ must be made clear here. Because of different theta functions, the divergent part and finite part have different boundaries. For divergent part, the theta function is $\theta(x_{2}-x^{*}_{2})$, so, $x_{2}>x^{*}_{2}=\xi>0$. On the other hand, for $T_{F}(x_{1},x_{2})$, the support is $|x_{1}|<1$, $|x_{2}|<1$ and $|x_{1}-x_{2}|<1$, so, for $T_{F}(x^{*}_{2},x^{*}_{2}-x_{2})$, we have $|x_{2}|<1$. Together with $x_{2}>\xi$, we have $\xi<x_{2}<1$. Keeping the divergent part only, we have
$I\langle L\rangle\Big{|}_{v}\doteq$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta 2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big{]}$
$+\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})g_{1}^{(c)}\Big{\}},$ (168)
where $\int_{r}$ is defined by eq.(120). Without $g_{1}^{(c)}$, above result is 4-4 times of $I\langle P_{7}\rangle|_{v}$.
The real corrections of $I\langle L\rangle$ are given by eqs.(123, 128,131,134). The divergent part is summarized as follows:
$I\langle L\rangle\Big{|}_{r}\doteq$ $2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(a)}+\delta(1-\hat{x}_{b})g_{1}^{(a)}+\delta(1-\hat{x}_{2})g_{2}^{(a)}\Big{]}$
$+$ $\bar{q}(x_{b})T_{F}(x^{*}_{1},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(b)}+\delta(1-\hat{x}_{b})g_{1}^{(b)}+\delta(1-\hat{x}_{2})g_{2}^{(b)}\Big{]}$
$+$ $\bar{q}(x_{b})T_{F}(x^{*}_{1},x^{*}_{1}-x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(c)}+\delta(1-\hat{x}_{b})g_{1}^{(c)}+\delta(1-\hat{x}_{2})g_{2}^{(c)}\Big{]}\Big{\}},$ (169)
with
$g_{0}^{(a)}=$ $-\frac{256}{\epsilon^{2}},$
$g_{0}^{(b)}=$ $\frac{256N_{c}^{2}}{\epsilon^{2}},$
$g_{0}^{(c)}=$ $0;$ (170)
and
$g_{1}^{(a)}=$ $\frac{64(\hat{x}_{2}^{2}+1)}{\epsilon(1-\hat{x}_{2})_{+}},$
$g_{1}^{(b)}=$ $-\frac{64N_{c}^{2}(1+\hat{x}_{2})}{\epsilon(1-\hat{x}_{2})_{+}},$
$g_{1}^{(c)}=$ $-\frac{64}{\epsilon};$ (171)
and
$g_{2}^{(a)}=$ $\frac{64(\hat{x}_{b}^{3}+\hat{x}_{b})}{\epsilon(1-\hat{x}_{b})_{+}},$
$g_{2}^{(b)}=$ $-\frac{64(1+\hat{x}_{b}^{2})(\hat{x}_{b}+N_{c}^{2}-1)}{\epsilon(1-\hat{x}_{b})_{+}},$
$g_{2}^{(c)}=$ $0.$ (172)
Now, the sum of virtual and real corrections give
$g_{0}|_{r+v}=$ $g_{0}|_{v}+g_{0}|_{r}^{a+b+c}=-\frac{192(N_{c}^{2}-1)}{\epsilon}+32(N_{c}^{2}-1)(\pi^{2}-8).$ (173)
However, the subtraction term is
$g_{0}|_{sub}=$ $-\frac{-192+64N_{c}^{2}}{\epsilon}+32(N_{c}^{2}-1)(\pi^{2}-8).$ (174)
So,
$g_{0}|_{r+v}-g_{0}|_{sub}=-\frac{128N_{c}^{2}}{\epsilon}.$ (175)
This is the divergence that cannot be subtracted!
Another problem is for the non-pole virtual correction. In the above we have written this part into $g_{1}^{(c)}$, i.e.,
$I\langle L\rangle\Big{|}_{v}\supset$ $\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\theta(x_{2}-x_{2}^{*})g_{1}^{(c)},\ g_{1}^{(c)}=128\hat{x}_{2}(\frac{1}{\epsilon}+1),$
$\text{const.}=$ $(l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta)2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}.$ (176)
We notice that when $\hat{x}_{b}=1$, $x_{2}^{*}=x_{1}^{*}$ with $x_{1}^{*}$ determined by the hard pole condition $(k_{1}-q)^{2}=0$. When $\hat{x}_{b}=1$, the magnitude of the transverse momentum of virtual photon is
$q_{t}=\sqrt{\frac{Q^{2}}{\hat{x}_{b}}(1-\hat{x}_{2})(1-\hat{x}_{b})}=0.$ (177)
Then,
$q^{-}=\frac{Q^{2}+q_{t}^{2}}{2q^{+}}=\frac{Q^{2}}{2\xi p_{a}^{+}}.$ (178)
So,
$x_{1}^{*}=\frac{Q^{2}}{2p_{a}\cdot q}=\xi.$ (179)
This is just $x_{2}^{*}$. Thus, non-pole contribution can be written as
$I\langle L\rangle\Big{|}_{v}\supset\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})\delta(1-\hat{x}_{b})g_{1}^{(c)},$ (180)
and is actually a part of hard pole contribution. After adding it to corresponding real contribution, we find
$I\langle L\rangle\Big{|}_{v}+I\langle L\rangle\Big{|}_{r}\supset$ $\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})\delta(1-\hat{x}_{b})\Big{[}-4\frac{16-32\hat{x}_{2}}{\epsilon}+128\hat{x}_{2}\Big{]}.$ (181)
The divergence is the same as that of subtraction term,
$I\langle L\rangle\Big{|}_{sub}\supset-4\times\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\frac{16-32\hat{x}_{2}}{\epsilon}.$ (182)
Thus, the divergent non-pole part in virtual correction is safe, and should be viewed as a part of hard pole contribution.
Other $g_{i}$’s are also safe:
  1. 1.
    $g_{1}^{a}$:
    $g_{1}^{(a)}|_{r+v}=g_{1}|_{r}^{(a)}=\frac{64(1+\hat{x}_{2}^{2})}{\epsilon(1-\hat{x}_{2})_{+}},\ g_{1}^{(a)}|_{sub}=-4\times\frac{-16(1+\hat{x}_{2}^{2})}{\epsilon(1-\hat{x}_{2})_{+}},$ (183)
    so,
    $g_{1}^{(a)}|_{r+v}-g_{1}^{(a)}|_{sub}=0.$ (184)
  2. 2.
    $g_{1}^{(b)}$:
    $g_{1}|_{r+v}^{(b)}=g_{1}|_{r}^{(b)}=-\frac{64N_{c}^{2}(1+\hat{x}_{2})}{\epsilon(1-\hat{x}_{2})_{+}},\ g_{1}|_{sub}^{(b)}=-4\times\frac{16N_{c}^{2}(1+\hat{x}_{2})}{\epsilon(1-\hat{x}_{2})_{+}}.$ (185)
    So,
    $g_{1}|_{r+v}^{(b)}-\ g_{1}|_{sub}^{(b)}=0.$ (186)
  3. 3.
    $g_{2}$:
    $g_{2}|_{r+v}=g_{2}|_{r}^{(a+b+c)}=-\frac{64(N_{c}^{2}-1)(1+\hat{x}_{b}^{2})}{\epsilon(1-\hat{x}_{b})_{+}},\ g_{2}|_{sub}=-4\times\frac{16(N_{c}^{2}-1)(1+\hat{x}_{b}^{2})}{\epsilon(1-\hat{x}_{b})_{+}}.$ (187)
    So,
    $g_{2}|_{r+v}-g_{2}|_{sub}=0.$ (188)
In summary, all divergences can be removed by collinear subtraction, except for $g_{0}$-term, i.e.,
$I\langle L\rangle\Big{|}_{r+v}-I\langle L\rangle\Big{|}_{sub}=2\times\frac{\alpha_{s}l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})\frac{-128N_{c}^{2}}{\epsilon}\Big{]}+\text{finite.}$ (189)
This indicates twist-3 collinear factorization does not hold. In $\cdots$], $N_{c}^{2}$ should be understood as $C_{A}N_{c}$, and this term comes from Fig.5(c). As an example, we present the calculation of HP contribution of Fig.5(c) in App.B. More clearly, we redraw the diagrams in Fig.8. We analyze the divergence of these two diagrams in next section with eikonal approximation.
VIII Eikonal approximation for HP contribution
From the result of last section, it is the divergence from HP breaks the factorization. The divergence is proportional to $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$, which indicates the final gluon with momentum $k_{g}$ is a soft gluon. The two diagrams having HP contribution proportional to $C_{A}N_{c}$ are shown in Fig.8.
Figure 8: The diagrams having HP contribution proportional to $C_{A}N_{c}$.
The contribution proportional to $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$ can be given by soft approximation or eikonal approximation. The final gluon in this approximation is a soft one, with $k_{g}^{\pm}\sim k_{g\perp}\sim\lambda$. $\lambda$ is a small quantity. The hard pole condition is $(k_{1}-q)^{2}=0$ or $(k_{b}+k-k_{g})^{2}=0$. So,
$2k_{b}^{-}(k^{+}-k_{g}^{+})+k_{g\perp}^{2}=0.$ (190)
Because $k_{g\perp}\sim\lambda$, we must have
$k^{+}-k_{g}^{+}\sim\lambda^{2},$ (191)
which also implies $k^{+}\sim\lambda$. This is different from twist-2 case, where all external momenta are $O(1)$.
To simplify the analysis, we use $G_{\perp}$ as the initial gluon from polarized hadron to do collinear expansion. Then,
$I\langle L\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})L_{\mu\nu}W^{\mu\nu}$
$=$ $\int\frac{dq^{+}}{2q^{+}}\int d^{n-2}q_{\perp}\int dk_{2}^{+}dk^{+}dk_{b}^{-}\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}\tilde{s}_{\perp}^{\rho}\frac{i}{k^{+}}(-i\pi)\delta((k_{1}-q)^{2})\hat{H}_{\rho}^{\mu\nu}L_{\mu\nu},$ (192)
where
$\hat{H}_{\rho}^{\mu\nu}=$ $\int\frac{d^{n}k_{g}}{(2\pi)^{n}}2\pi\delta_{+}(k_{g}^{2})\delta^{n}(k_{2}+k_{b}-q-k_{g})H_{1,2}^{\mu\nu,\rho}$
$=$ $\frac{1}{(2\pi)^{n-1}}\hat{x}_{2}\delta(q_{t}^{2}-\frac{Q^{2}}{\hat{x}_{2}}(1-\hat{x}_{2})(1-\hat{x}_{b}))H_{1,2}^{\mu\nu,\rho}.$ (193)
$H_{1,2}^{\mu\nu,\rho}$ is the hard part for Fig.8(a) and (b), respectively. The explicit expressions are
[…]
with $\Gamma_{\rho\tau\lambda}(k,k_{g}-k,-k_{g})=g_{\rho\tau}(2k-k_{g})_{\lambda}+g_{\tau\lambda}(2k_{g}-k)_{\rho}+g_{\lambda\rho}(-k_{g}-k)_{\tau}$. After integrating out $q_{\perp}$ and $k^{+}$, we get
$I\langle L\rangle=$ $\frac{\pi\Omega_{n-2}}{4(2\pi)^{n-1}}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}\tilde{s}_{\perp\rho}(q_{t}^{2})^{-\epsilon/2}\frac{1}{2k^{+}q^{-}}H_{1,2}^{\mu\nu,\rho}L_{\mu\nu},\ q_{t}^{2}=\frac{Q^{2}}{\hat{x}_{2}}(1-\hat{x}_{2})(1-\hat{x}_{b}).$ (195)
The leading power behavior of the integrand is
$\frac{1}{2k^{+}q^{-}}H_{1,2}^{\mu\nu,\rho}L_{\mu\nu}\sim\frac{1}{q_{\perp}^{2}}\sim\frac{1}{\lambda^{2}}.$ (196)
After expansion in $\epsilon$, this part gives $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$. Keeping leading power contribution only, we get
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\tilde{s}_{\perp}^{\rho}(q_{t}^{2})^{-\epsilon/2}\Big{[}\frac{N_{1}^{\rho}}{2k\cdot q[(k_{b}-k_{g})^{2}][(k_{g}-k)^{2}]}+\frac{N_{2}^{\rho}\Big{|}_{\text{II}}+N_{2}^{\rho}\Big{|}_{\text{III}}}{2k\cdot q[(k-k_{g})^{2}][(k_{2}-k_{g})^{2}]}\Big{]},$
$\text{K.}=$ $\frac{\Omega_{n-2}}{4(2\pi)^{n-1}}\frac{[-ig_{s}^{2}f^{abc}Tr(T^{c}T^{b}T^{a})]}{8N_{c}^{2}C_{F}}.$ (197)
Because $^{+}\sim\lambda, k+k$$^{+}-k_{g}^{+}\sim\lambda^{2}$, the two denominators are of order $\lambda^{4}$. $N_{1}^{\rho}$ is given by Fig.8(a) and $N_{2}^{\rho}|_{\text{II}}$, $N_{2}^{\rho}|_{\text{III}}$ are given by Fig.8(b). Their expressions are
$N_{1}^{\rho}=$ $-16k_{g}\cdot k_{b}[k^{\nu}g_{\perp}^{\mu\rho}+k^{\mu}g_{\perp}^{\nu\rho}]L_{\mu\nu},$
$N_{2}^{\rho}|_{\text{II}}=$ $\frac{32q_{\perp}^{2}}{n-2}L_{\mu\nu}\Big{[}-\Big{(}\frac{\partial}{\partial q_{\perp}}\Big{)}_{L}g_{\perp}^{\mu\nu}+[k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho}]+\frac{1}{2}\Big{(}[k_{b}^{\nu}g_{\perp}^{\mu\rho}+k_{b}^{\mu}g_{\perp}^{\nu\rho}]-[k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho}]\Big{)}\Big{]},$
$N_{2}^{\rho}|_{\text{III}}=$ $16k\cdot k_{g}[k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho}]L_{\mu\nu}.$ (198)
All of them are of order $\lambda^{2}$. In the above, $L^{\mu\nu}$ has been expanded in $q_{\perp}$ at $q_{\perp}=0$. The derivative $(\partial/\partial q_{\perp}^{\rho})_{L}$ in $N_{2}^{\rho}$ indicates the derivative acts on $L^{\mu\nu}$. After taking the derivative, $q_{\perp}$ is set to zero. The benefit of this decomposition is $N_{2}^{\rho}|_{\text{II}}$ alone satisfies QED gauge invariance, while $N_{1}^{\rho}$ and $N_{2}^{\rho}|_{\text{III}}$ together recovers such a symmetry. As can be seen,
$\frac{N_{1}^{\rho}}{2k\cdot q[(k_{b}-k_{g})^{2}][(k_{g}-k)^{2}]}+\frac{N_{2}^{\rho}\Big{|}_{\text{III}}}{2k\cdot q[(k-k_{g})^{2}][(k_{2}-k_{g})^{2}]}=\frac{4L_{\mu\nu}}{k_{2}\cdot qq_{\perp}^{2}}\Big{[}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{]}.$ (199)
If we replace $L^{\mu\nu}$ by $q^{\mu}$ or $q^{\nu}$, the above is zero. This tensor also appears in $N_{2}|_{\text{II}}$. Then, the sum is
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\tilde{s}_{\perp}^{\rho}(q_{t}^{2})^{-\epsilon/2}\frac{4L_{\mu\nu}}{k_{2}\cdot qq_{\perp}^{2}}\Big{\{}\frac{4}{2-\epsilon}\Big{[}-k_{2}\cdot k_{b}g_{\perp}^{\mu\nu}\Big{(}\frac{\partial}{\partial q_{\perp}^{\rho}}\Big{)}_{L}+(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})\Big{]}$
$+(1-\frac{2}{2-\epsilon})\Big{[}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{]}\Big{\}}.$ (200)
Because
$g_{\perp\mu\nu}\Big{(}\frac{\partial L^{\mu\nu}}{\partial q_{\perp}^{\rho}}\Big{)}_{q_{\perp}=0}=0,$
$L_{\mu\nu}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})\Big{|}_{q_{\perp}=0}=-L_{\mu\nu}(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{|}_{q_{\perp}=0}=8k_{2}\cdot k_{b}\cos\theta l_{\perp,cs}^{\rho},$ (201)
we have
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})(q_{t}^{2})^{-\epsilon/2}\frac{64\cos\theta l_{\perp,cs}\cdot\tilde{s}_{\perp}}{q_{\perp}^{2}}\Big{\{}\frac{2}{2-\epsilon}+(1-\frac{2}{2-\epsilon})\Big{\}}.$ (202)
Now the two $2/(2-\epsilon)$ cancel out. From above derivation we see clearly how the cancellation occurs. The divergence of $I\langle L\rangle$ is given by
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})64\cos\theta l_{\perp,cs}\cdot\tilde{s}_{\perp}\frac{(q_{t}^{2})^{-\epsilon/2}}{q_{\perp}^{2}}$
$=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\frac{-64\cos\theta l_{\perp,cs}\cdot\tilde{s}_{\perp}}{Q^{2}}\Big{(}\frac{4}{\epsilon^{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})+\cdots\Big{)}.$ (203)
where $\cdots$ in brackets does not contain the two delta functions at the same time.
Above formulas for $I\langle L\rangle$ also hold for $I\langle P_{7}\rangle$. But because $P_{7}\propto q_{\perp}$, only the derivative term in $N_{2}^{\rho}|_{\text{II}}$ contributes. Because
$g_{\perp\mu\nu}\frac{\partial P_{7}^{\mu\nu}}{\partial q_{\perp}^{\rho}}=\frac{1}{\tilde{s}_{\perp}^{2}}g_{\perp\mu\nu}(g_{\perp}^{\mu\rho}\tilde{s}_{\perp}^{\nu}+g_{\perp}^{\nu\rho}\tilde{s}_{\perp}^{\mu})=\frac{2\tilde{s}_{\perp}^{\rho}}{\tilde{s}_{\perp}^{2}},$ (204)
we have
$I\langle P_{7}\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\frac{(q_{t}^{2})^{-\epsilon/2}}{q_{\perp}^{2}}\frac{-32}{2-\epsilon}.$ (205)
The factor $1/(2-\epsilon)$ is left. Now, it is clear that under eikonal approximation, the divergent part of $I\langle L\rangle$ and $I\langle P_{7}\rangle$ satisfies following relation
$I\langle L\rangle=-4(1-\frac{\epsilon}{2})\tilde{s}_{\perp}\cdot l_{\perp,cs}\cos\theta I\langle P_{7}\rangle,$ (206)
which is consistent with our previous explicit results eqs.(127,128).
Moreover, if there is no the second line in eq.(200), we have $I\langle L\rangle=-4I\langle P_{7}\rangle$, then the factorization holds. The tensor in the second line of eq.(200),
$(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})$ (207)
has a coefficient proportional to
$(1-\frac{2}{2-\epsilon})=\frac{-\epsilon}{2-\epsilon}.$ (208)
But because $1/q_{\perp}^{2}$ gives $1/\epsilon^{2}$, this tensor still has a divergent coefficient. It is this coefficient that cannot be subtracted. The same tensor also appears in virtual correction. See eq.(91), where
$g_{3,NP}^{\mu\nu}=C_{3,NP}t_{a}^{\mu\nu}+D_{3,NP}t_{b}^{\mu\nu}=C_{3,NP}\frac{1}{x_{2}}\Big{[}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{]}.$ (209)
However, $g_{3,NP}^{\mu\nu}$ has only non-pole contribution($x_{1}\neq x_{2}$), which cannot cancel the divergence given here. Thus, the soft divergence associated with above tensor in real correction cannot be cancelled by corresponding virtual correction. Moreover, from eq.(191), the gluon with momentum $k_{g}-k$ behaviors like $_{g}^{+}-k^{+}\sim\lambda^{2}, kk$$_{\perp}-k_{g\perp}\sim\lambda$($k_{\perp}=0$). So, $|(k_{g}^{+}-k^{+})(k_{g}^{-}-k^{-})|\ll|k_{g\perp}-k_{\perp}|^{2}$. The gluon thus is a Glauber gluon. In unpolarized DY, the Glauber gluon contribution is shown to cancel out, see for example Bodwin (1985). Our above analysis indicates that such a cancellation does not happen for SSA.
IX Summary
In this work, we have calculated the one-loop correction to unweighted SSA in DY for the lepton angular distribution. As a comparison, we also calculate a weighted observable $I\langle P_{7}\rangle$ with the same method. We use at most one $G^{+}$ to do collinear expansion, and then use EOM for fermion to eliminate the bad component of fermion field. After this, the hadronic tensor or cross section is expressed by independent twist-3 distribution functions. Three of these functions are $q_{\partial}$, $T_{F}$ and $T_{\Delta}$. For virtual correction, we first determine that the hard part is analytic on the upper half planes of $s_{0},s_{1}$. After this is clear, we use FIRE to reduce the various tensor integrals in the hard part to standard scalar integrals. Due to the simple kinematics, the reduced scalar integrals are three bubble integrals $B(s_{0})$, $B(s_{1})$ and $B(s_{2})$. Then, the structure of virtual correction becomes clear. We successfully extract pole part and non-pole part. For pole part, only SGP is possible. QCD and QED gauge invariance are maintained in the calculation. On the other hand, for real correction, we first assume $q_{\perp}\neq 0$. Then, the tensor structure of $W^{\mu\nu}$ is worked out. With the projection tensors we calculate the real corrections to $I\langle L\rangle$ in the same way as $I\langle P_{7}\rangle$. Only three-point distribution functions contribute at one-loop level. The contributions are classified into SGP, HP and SFP contributions. Corresponding contributions are worked out explicitly. In this work, only contributions from $\bar{q}(x_{b})$ and $T_{F}(x_{1},x_{2})$ are retained. For both virtual and real corrections, $I\langle P_{7}\rangle$ is the same as our previous resultChen et al. (2017). Then, the collinear subtraction is performed. We find all divergences in $I\langle P_{7}\rangle$ can be subtracted out, while a single pole in $I\langle L\rangle$ cannot be subtracted. The single pole is proportional to $N_{c}C_{A}/\epsilon$ and $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$. In addition, we find the divergent nonpole part of virtual correction can be viewed as a part of HP contribution of real correction, and can be subtracted without any problem. The only term to break the factorization is the single pole term we mentioned above. We also analyze the source for this unsubtracted divergence under eikonal approximation for real HP corrections. We find in $W^{\mu\nu}$ a tensor appears with a soft divergence, but such a tensor in virtual correction has no corresponding soft divergence. We thus conclude that it is this uncancelled soft divergence that breaks the factorization. Very recently, the one-loop correction to SSA in lepton-hadron scattering is calculated in Rein et al. (2025a, b). No breaking of factorization is found. In future, we will study this process using present method. Moreover, the breaking of factorization found here depends on collinear expansion. We would like to check this conclusion in future with the multiparton state proposed in Cao et al. (2010); Ma et al. (2012).
Acknowledgements
The author would like to thank J.P. Ma for very helpful discussions. The hospitality of ITP, CAS(Beijing) is appreciated during the finalization of this paper.
Appendix A Counter terms of QCD
The counter terms we need in this work are
[…]
All fields and coupling constants are renormalized ones. $G^{\mu}=G_{a}^{\mu}T^{a}$ is gluon field, and $A^{\mu}$ is photon field. The relation between bare and renormalized quantities is
$\psi_{B}=Z_{2}^{1/2}\psi,\ G_{B,a}^{\mu}=Z_{3}^{1/2}G_{a}^{\mu},\ g_{s}^{B}=\frac{Z_{1F}}{Z_{2}Z_{3}^{1/2}}g_{s}.$ (211)
In $\overline{MS}$ scheme, the values of these constants are well known, and can be found in e.g.,Muta (1998). For convenience, we list the values in Feynman gauge as follows:
$\delta z_{3}=$ $Z_{3}-1=-\frac{g_{s}^{2}}{16\pi^{2}}R_{\epsilon}(\frac{2}{3}N_{F}-\frac{5}{3}C_{A}),$
$\delta z_{2}=$ $Z_{2}-1=-\frac{g_{s}^{2}}{16\pi^{2}}R_{\epsilon}C_{F},$
$\delta z_{1F}=$ $Z_{1F}-1=-\frac{g_{s}^{2}}{16\pi^{2}}R_{\epsilon}(C_{A}+C_{F}),$
$\delta z_{1}^{\gamma}=$ $Z_{1}^{\gamma}-1=-\frac{g_{s}^{2}}{16\pi^{2}}R_{\epsilon}C_{F},$ (212)
with $R_{\epsilon}=\frac{2}{\epsilon}-\gamma_{E}+\ln 4\pi$. $N_{F}$ is the number of quark flavor. Note that $Z_{1}^{\gamma}=Z_{2}$.
Appendix B Calculation of Fig.5(c) with transverse gluon
Since Fig.5(c) causes the problem, we give the details for this diagram. For hard pole, $k^{+}\neq 0$, we can use $G_{\perp}$ to calculate directly. With $G_{\perp}$ as the initial gluon, the hard part is
$H^{\mu\nu}_{\rho}=$ $\int\frac{d^{n}k_{g}}{(2\pi)^{n}}2\pi\delta_{+}(k_{g}^{2})\delta^{n}(k_{2}+k_{b}-q-k_{g})$
[…]
$\frac{-i}{(k_{g}-k)^{2}+i\epsilon}\pi\delta((k_{1}-q)^{2}).$ (213)
Then, we integrate out $k_{g}$ using the delta function for momentum conservation, and get
$H^{\mu\nu}_{\rho}=$ $\frac{\pi}{(2\pi)^{n-1}}\delta_{+}((k_{2}+k_{b}-q)^{2})$
[…]
$\frac{-i}{(k_{g}-k)^{2}+i\epsilon}\delta((k_{1}-q)^{2})\Big{|}_{k_{g}\rightarrow k_{2}+k_{b}-q}.$ (214)
For $G_{\perp}$ expansion, partons have no transverse momentum, so,
$k_{2}^{\mu}=x_{2}p_{a}^{\mu},\ k_{1}^{\mu}=x_{1}p_{a}^{\mu},\ k^{\mu}=xp_{a}^{\mu},\ k_{b}^{\mu}=x_{b}p_{b}^{\mu}.$ (215)
So, in $H^{\mu\nu}_{\rho}$ there is only one transverse momentum $q_{\perp}^{\mu}$. Then,
$I\langle L\rangle=\int d^{n}q\delta(q^{2}-Q^{2})\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}dk_{1}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}\tilde{s}_{\perp}^{\rho}\frac{i}{k^{+}}L_{\mu\nu}H^{\mu\nu}_{\rho}$ (216)
Then, we integrate out $k_{1}^{+}$ with the pole condition $\delta((k_{1}-q)^{2})$, which gives $x_{1}=x_{1}^{*}=Q^{2}/(2p_{a}\cdot q)$. In $L_{\mu\nu}$, $l^{\mu}$ is given by eq.(8). This introduces another transverse momentum $l_{\perp,cs}^{\mu}$. In the simplifying of the result, we use following replacement
$l_{\perp,cs}\cdot q_{\perp}q_{\perp}^{\rho}\rightarrow\frac{l_{\perp,cs}^{\rho}q_{\perp}^{2}}{n-2}.$ (217)
This is due to the symmetry of $q_{\perp}$ integration. After this, we get a result with integrand depending on $q_{\perp}^{2}$ only, that is,
$I\langle L\rangle=$ $\tilde{s}_{\perp}^{\rho}\int d^{n}q\delta(q^{2}-Q^{2})\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\delta((k_{2}+k_{b}-q)^{2}).$ (218)
Then, we integrate out $q^{-}$ using $\delta(q^{2}-Q^{2})$ and then let $q^{+}=\xi p_{a}^{+}$. $\delta((k_{2}+k_{b}-q)^{2})$ can be simplified as
$\delta((k_{2}+k_{b}-q)^{2})=\hat{x}_{2}\delta(\frac{Q^{2}}{\hat{x}_{b}}(1-\hat{x}_{2})(1-\hat{x}_{b})-q_{t}^{2}).$ (219)
This enables us to integrate out $q_{\perp}$ and get
$I\langle L\rangle=$ $\tilde{s}_{\perp}^{\rho}\int\frac{d\xi}{2\xi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\int d^{n-2}q_{\perp}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\delta((k_{2}+k_{b}-q)^{2})$
$=$ $\tilde{s}_{\perp}^{\rho}\int\frac{d\xi}{2\xi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\frac{\Omega_{n-2}}{2}(q_{t}^{2})^{-\epsilon/2}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\hat{x}_{2}.$ (220)
Note that in $H^{\mu\nu}_{\rho}$, all Lorentz vectors are defined in n-dim space, including $\rho$. $L_{\mu\nu}$ is also defined in n-dim space. So, $l_{\perp,cs}^{\rho}$ is in n-dim space. After contracting with $\tilde{s}_{\perp}$, it falls into 4-dim space. Anyhow, this does not affect the calculation of $\bar{H}$. With some simplifications, we have
$I\langle L\rangle=$ $\tilde{s}_{\perp}\cdot l_{\perp,cs}\frac{\pi\Omega_{n-2}}{4(2\pi)^{n-1}}\int\frac{d\xi}{\xi}\int\frac{x_{b}}{x_{b}}\bar{q}(x_{b})\int\frac{dx_{2}}{x_{2}}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}(q_{t}^{2})^{-\epsilon/2}[\cdots],$ (221)
with
$[\cdots]=$ $\frac{g_{s}^{2}C_{A}}{16N_{c}}\frac{32\cos\theta\hat{x}_{b}}{(2-\epsilon)Q^{2}(1-\hat{x}_{2})(1-\hat{x}_{2})}\Big{[}(1-\epsilon)E_{t}Q\hat{x}_{2}+Q^{2}[1+\hat{x}_{2}(\hat{x}_{b}-1)][1+(2-\epsilon)\hat{x}_{2}(\hat{x}_{b}-1)]\Big{]}.$ (222)
A nontrivial feature is the hard part is automatically proportional to $\cos\theta$. Writing the expression into standard form, we have
$I\langle L\rangle=$ $\tilde{s}_{\perp}\cdot l_{\perp,cs}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\Big{\{}C_{A}N_{c}\frac{64\hat{x}_{b}}{(2-\epsilon)Q^{2}(1-\hat{x}_{2})(1-\hat{x}_{b})}\Big{(}\frac{Q^{2}}{q_{t}^{2}}\Big{)}^{\epsilon/2}[\cdots]\Big{\}}.$ (223)
Since there are singularities at $\hat{x}_{2}=1$ and $\hat{x}_{b}=1$, the expansion in $\epsilon$ is realized by
$\frac{1}{(1-x)^{1+\epsilon/2}}=-\frac{2}{\epsilon}\delta(1-x)+\frac{1}{(1-x)_{+}}-\frac{\epsilon}{2}\Big{(}\frac{\ln(1-x)}{1-x}\Big{)}_{+}.$ (224)
Then, we get
$\Big{\{}\cdots\Big{\}}=$ $C_{A}N_{c}\frac{64}{2-\epsilon}\frac{4}{\epsilon^{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})[(1-\epsilon)+1]+\cdots$
$=$ $C_{A}N_{c}\frac{256}{\epsilon^{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})+\cdots.$ (225)
$\cdots$ represents other terms which do not contain two delta functions. This is the result given in eq.(128). As can be seen, $(2-\epsilon)$ in denominator is cancelled by the same factor in numerator. So, only $1/\epsilon^{2}$ is left.